A Bibliography of Publications of Sir Maurice V. Wilkes

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

17 April 2025 Version 1.10

Abstract

This bibliography records publications of the late Sir Maurice V. Wilkes.

Title word cross-reference

#3 [Gor72].

(1, 1) [LMW79]. Ch(x, χ) [Wil54].

/etc/init.d/halt [Orl10].

1 [Har99, O’R18c, WM72]. 1-85196-040-6 [Wil95b]. 10th [WK51]. 11th [WK51]. 12th [WK51]. 1940s [Ley11]. 1951 [McP52]. 1957 [GOB^{+}57, Cor58]. 1961 [Wil64f]. 1973 [GLL74]. 1977 [Gil77]. 1978 [Wil78e]. 1980s [Wil80a, Wil80c]. 1984 [Ano85]. 1996 [BS96]. 1997 [Wil97c].

2 [WW58, Wil92h]. 2001 [WK03]. 2003 [HW04]. 2010 [CK14]. 2013 [Sco14]. 2014 [KT14]. 2015 [Ste16]. 2016 [Mar17]. 2018 [Dwa18]. 24th [BS96]. 28a [Wil00e]. 2nd [Cor58].

342 [Wil96d]. 360 [Tuc67]. 3rd [GOB^{+}57, Wil78e].

40 [Wil74d]. 40th [MHW87].

6.75 [Wil72b]. 60 [SW61, Wil76b].

77 [Wil03]. ’78 [Wil78e]. 7th [Gil77].

8th [GOB^{+}57].

91 [AEGP67]. 93 [Ano88]. ’96 [BS96]. 9th [WK51].

Aberdeen [Ber00]. Abstracts [Wil65a]. Abuse [Wil03]. Academic [Wil72b].

Access [Wil67b, Wil68a, Wil73a, Wil64b, Wil67e, WN68]. according [Wil40b]. Account [Wil98c]. ACE [AWL^{+}88, Cop12]. ACM [BS96, Ash87]. act [BW00b, BW00a, Wil00d]. Action [Wil64f]. Activity [CK80b, CK92c]. Addison [Cor58]. Addison-Wesley [Cor58]. addres [Wil53g]. address [Wil82c]. Adjustment [Wil56c, Wil71b, Wil71c]. Advances [Wil72b]. Aerial [Bly57]. after [Har99]. Age [Cur12a]. aged [Ano10]. Ages [Kid99]. Ahead [SG69]. AIEE [McP52]. AIEE-IRE [McP52]. Aiken [BHW^{+}90, CWC99]. Airy [CK92a]. AL [WM72]. AL-1 [WM72]. Alan [Ben12, Cop05, Cop12]. Algebraic [Bro53, Wil66a, Wil56d]. ALGOL [SW61, Wil93a]. algorithm [Wil56d]. Allan [Wil00e]. Allies [AWL^{+}88]. Allocation [Wil65e, Wil69c, Wil00c]. allowance [Wil63b]. Alt [Wil72b]. Amplifiers [Wil93b]. Analog [Bru20]. Analyse [Wil67a]. Analyser [Wil76d, LJWB39]. Analysis [SG69, Wil65b, Wil66b, Wil67a, Ken63, Bre68]. Analytical [Wil00e]. Anecdotes [Dou79, TW87, vdPWB82]. Anniversary [MHW87]. Annual [BS96]. Anticipation [BHW^{+}90]. Appleton [Wil97d]. Application [Gil53, Wil77a, Wil04a, GHR57, Wil77b]. Applications [Ben50, HJ04, Whe92b]. applied [Wil53f]. Approach [Wil84e, Wil92a]. Approaches [Wil92e]. approximations [LMW79]. April [IEE92]. Architects [Cur12a, Kid99]. Architecture [GL03, HJS00, KT14, Wil72c, Ken63, Wil80f]. Architectures [RH00, Wil78b]. Archives [Ano06]. Area [Les83, Wil80e]. Areas [Wil04a]. argument [SW50]. Arithmetic [Wil97a]. Army [Ber00]. Art [Wil74a, Wil74d, Wil76b]. Artificial [AMW^{+}88, Wil92e]. Aspects [Tay52, Lip68, Wil59c]. Aspray [Wil96d]. Association [Wil90c]. Associative [Wil71a, Wil72a, WM72]. Astronomy [Bly57]. Atanasoff [AMW^{+}88, BHW^{+}90]. Atlanta [Wil78e]. Atmosphere [Wil49c, Wil63b, Wil52b]. Atmospheric [WW47, Wil49a, Wil51d]. attachment [Wil40b]. Aug [Wil03]. August [Gil77, WK03]. Automatic [AWL^{+}88, BS54, BW53b, GOB^{+}57, Max57, Whe51, Wil50a, WR50b, Wil51a, WPB54, WWG54, Wil56a, Wil57a, Wil59a, Wil71b, Wil71c, Wil76a, Wil96a, Gil51b, Wil49d, Wil51b, Wil53g, Wil81a, Wil89a, WPB89, Wil21, Cop05, Wil60]. Automativ [Wil56c]. Auxiliary [Wil57c, WW58, WW56a, WW56b]. Avtomaticeskie [Wil60]. Award [Ano85, Ash87, KT14, Ano90, Ano91, Ano08, Dwa18, HT81, KT14, Mar17, Sco14, Ste16].

Babbage [BHW^{+}90, Wil64f, Wil95b, Wil00e, BWV^{+}88, BHW^{+}90, BW00b, BW00a, TW87, Wil64f, Wil71d, Wil75g, Wil76c, Wil77c, Wil81b, Wil83a, Wil90i, Wil91b, Wil91c, Wil91d, Wil91h, Wil92d, Wil92f, Wil95b, Wil00a, Wil00d, Wil02b]. Baby [O’R18c]. Back [Den99, WS86]. Bakerian [Ryl58]. band [Wil80e]. bandwidth [Wil90b, WH97]. base [Wil74d, Wil76b]. based [Ley11]. Bases [Wil72d, Wil74c]. BASIC [WS86]. BasicBooks [Wil96d]. be [Wil99b]. Beam [Bly57]. Been [Wil91e]. before [Wil13, Wil90e]. Beginning [Gla97]. Behind [Cam03]. Berks [Wil74d]. Bernard [HW03]. Best [Ano08, Wil76a, Wil96a, Wil51b, Wil81a, Wil89a, Wil21]. Beyond [Wil77h]. big [Wil65d]. Bildgeschichte [dB05]. Biographies [HW03, BHW^{+}90]. Birthday [Lee93]. Biscayne [IEE92]. Bletchley [Ley11, Smi13]. Bloomfield [AMW^{+}88]. BNF [BHW^{+}90]. boffins [Smi13]. Book [Cor58, Hen86, Wil61a, Wil62a, Wil64f, Wil66a, Wil69a, Wil72b, Wil74d, Wil76b, Wil78a, Wil95b, Wil96d, Wil00a, Wil86c, Wil67a, Wil78b]. Boon [Wil74d]. Bowditch [BWV^{+}88, BHW^{+}90]. Brain [Cop12, Wil00a]. Brainerd [WG82]. Brief [Lav15]. Britain [CK80a, Cro90]. British [WCK89, Lav80b, Lav15, Orl10, Wil90c]. Broken [AWL^{+}88]. Bromley [Wil00e]. Buifendam [AWL^{+}88]. Build [Cop12, Cop05, Wil00a]. Built [Lav80b]. Bunyan [Wil74d]. Burks [AMW^{+}88, AMW^{+}88]. Business [Kid99, Wil78f, Wil79, Wil90d, Wil91f, Lan77, Lan00, Wil96e].

C [AWL^{+}88, Ber84, Wil74d]. cache [Lip68]. Calculating [Ano88, AWL^{+}88, WR49, Wil49e, Wil51a, Wil53a, WWG54, Wil64f, Wil76a, Wil96a, Wor53, BK52, Wil47a, Wil48, Wil49b, Wil49d, Wil89a, Wil91d, Wil21, Wil84b]. calculating-machine [Wil49b]. Calculator [Nau51a, WR50b, AWL^{+}88]. call [Nau51c]. Cambridge [BS82, CK80b, CK92b, CK92c, CK19a, Cro92, Lav80a, NH82, Nee92, PHL92, Wil49b, Wil53a, Wil73a, Wil75b, WN79, WW79, WN80, Wil80d, WU88, Wil90i, Wil04b]. came [Wil75g]. Campbell [WLC86, Wil95b, Wil96d]. Campbell-Kelly [WLC86, Wil95b, Wil96d]. Can [Wil51c, Wil53b, Wil53c, Wil57b, Wil98a, Wil99b]. Canada [Gil77]. CAP [Nee92, WN79]. capability [Wil80f, Wil82a, Wil82b]. Capsule [WS86, WLC86]. Carpenter [AWL^{+}88]. Carrées [Wil67a]. carried [Wil53e]. case [Lan00, Wil92a]. caught [BW00b, BW00a, Wil00d]. CBE [HW04, WK03]. Celebration [Ber00, MHW87]. Censuses [Wil61a]. Centre [vdPWB82]. Century [WS86, Wil99a, MHR80]. Certain [WW57, VRW^{+}45]. Challenges [Wil97b]. Changing [Wil71e, Wil94a]. Chapman [Wil54]. Chapter [CK92a]. character [Wil83a, Wil91h]. Charles [BHW^{+}90, Wil61a, Wil95b, Wil76c, Wil91d, Wil92d, Wil92f, Wil00a, Wil02b, Wil64f]. Chatto [Wil95b]. Checking [WPB54, WPB89]. Christos [Ste16]. Churchill [Wil97c]. cifrovye [Wil60]. Cipher [AWL^{+}88]. Circuits [WS71, Wil00b, Wil05, Wil13, WS53, WS82, Wil12]. Classic [Lew21]. Cline [Fis50]. CMOS [Wil95c, Wil96b, Wil98a]. Co [Cor58, Wil61a, Wil62a]. Cockcroft [Wil97c]. codebreaker [Cop05]. coefficient [LMW79]. Coefficients [BS54]. Cognitive [AWL^{+}88]. cogwheel [Wil00a]. Cohen [HW03]. collapsed [WH97]. Collected [AWL^{+}88]. Collection [MHR80]. College [WU88, Wil97c]. Colossus [BWV^{+}88, McL17, WG82]. comet [Nau51b]. Coming [Wil67b]. Commentaries [AWL^{+}88]. Comments [BWV^{+}88, BHW^{+}90, BW00b, Kis92, SW80, Wil00d]. Committee [vdPWB82, Wil90c]. Communication [Kid99, Wil75a, Wil80e]. communications [WW79]. Companion [O’R18a]. Compatible [CDD^{+}63]. Compendium [O’R18a]. compilation [Har99]. compiler [Wil63a, Wil64h]. Compiling [OSWW65a, OSWW65b, Wil63a, Wil64h]. complex [SW50]. Components [Wil62e]. Computation [AWL^{+}88, Ben50, BK52, BW53a, BS54, Nau51a, Wil50c]. Computer [Anoxx, AMW^{+}88, BS96, BN71, BSN82, BHK77, Bro59, CK80a, Cer13, CRB24, CWC99, Cop12, Cuf24, Eve52, Gil53, GL03, Hen86, HJ04, HJS00, Hop52, KT14, Kid99, Lap96, Lew21, MHW87, McP52, RM76, SW80, Swa76, WWG51, WWG57, WRW58a, Wil58c, Wil61b, Wil61d, Wil62b, Wil62c, Wil64d, Wil65a, Wil67c, Wil67d, Wil68a, Wil68c, Wil68d, Wil71d, Wil71e, WS71, Wil72c, Wil72e, Wil72g, Wil74a, Wil74b, Wil75b, Wil75f, Wil76e, Wil77f, Wil78b, WN79, WWG82, Wil84c, Wil85, WS86, WLC86, Wil90d, Wil91b, Wil91f, Wil91a, Wil86c, WCK89, Ano15, CD55, CD57, Cop05, Lan77, Lan00, Ley11, Oet52, Orl10, Smi13, Wil52a, WS53, Wil53g, WRW58b, Wil61c, Wil64b, Wil67e, WN68, Wil77c, WS82, Wil83b, Wil90h, Wil92b, Wil96e, Wil00a, WK51, Ano06, BWV^{+}88, Cor58, PL53]. Computer [WLC86, Ham58, WWG54, Wil62a, Wil96d]. Computers [CKW85, Lav80b, Lav15, McP52, O’R18b, O’R18c, Ran73, Ran75, Ran82, RH00, Str59b, Wil56a, Wil57a, Wil58a, Wil59a, Wil61e, Wil62d, Wil64e, Wil65f, Wil67b, Wil68e, Wil70, WR75, Wil77h, Wil77d, Wil78c, Wil80a, Wil80b, Wil80c, Wil87b, Wil88, Wil90e, Wil96c, Wil04a, Wil04b, Wil13, Wil56b, Wil60, Wil92a, BWV^{+}88, Max57, Wil72b]. Computing [Ber00, Bow53, BW53b, Bru20, CK92b, CK19a, Cer03, Cop05, Cop06, Cro90, Cro92, Eng50, Gor72, GOB^{+}57, GLL74, MHR80, NH82, O’R18a, Whe51, Wil50a, WW57, Wil78f, Wil79, Wil90f, Wil90g, Wil91e, Wil91f, Wil91g, Wil92e, Wil92f, Wil92g, Wil93a, Wil93b, Wil95a, dB05, Ano10, Gil51b, HNW^{+}48, Wil51b, Wil80e, Wil81a, Wil92d, Wil96b, dB05]. Conference [BS96, Bro59, GOB^{+}57, McP52, Wil59b, Wil78e, WK51]. Conferences [WCK89]. Congress [Gil77]. Conquest [AMW^{+}88]. Consequences [Wil91a]. Console [Wil64d]. Constable [Wil64f]. constant [LMW79]. Constraint [Wil64a]. Constraint-type [Wil64a]. construction [PL53]. Contribution [Wil97c]. Control [Wil58a, Wil62d, WS71, Wil72f, Wil81b, Tuc67, WS53, WRW58a, WRW58b, WS82]. Controlled [AWL^{+}88, Str59a]. Core [Wil69c]. Corrected [Cuf24]. Correspondence [HT81]. cosmic [Ryl58]. Could [Wil91e]. Coupling [BS54]. Created [Lew21]. Cryptologia [Kid99]. crystal [CD55, CD57]. Crystallographer [BW53a]. Crystallographic [Ken63]. CSC [BS96]. Current [Wil65a]. curriculum [Wil90i]. Curtain [Cam03]. curves [Wil39]. CXXIV [Oet52].

D [Ham58, Wil58b, Wil62a]. D. [Wil84b]. Data [GOB^{+}57, Wil61c, Wil72d, Wil74c, Wil78f, Wil79, Ken63, Wil71a, Wil72a, Wil74d, Wil76b]. data-processing [Ken63]. David [Cor58, Wil86c]. Davos [GLL74]. Days [Nau80]. deadly [Ley11]. Debate [BWV^{+}88, BHW^{+}90, Kis92, SW80]. Debates [BW00b, Wil00d]. Debugging [CK92a]. Decade [Wil58c]. December [Bro59, McP52]. Decisions [Wil94b, Wil90e]. Defense [Gro93]. Delay [Kis92, WR50b, Ano15]. delay-line [Ano15]. delivered [Wil04b]. Demonstration [Wor50]. Departments [Kid99]. Design [CKW85, Swa76, Wil48, Wil49e, Wil50b, Wil61d, WS71, Wil76a, Wil77a, Wil81b, Wil94b, Wil96a, LJWB39, Wil49d, Wil51b, WS53, WRW58b, WW59, Wil59c, Wil64b, Wil67e, WN68, Wil77b, Wil81a, WS82, Wil89a, Wil90e, Wil21]. designed [Wil86b]. Designers [Wil97b]. determination [CD55, CD57]. Determined [Fis50]. Deutsche [VRW^{+}45]. Development [AWL^{+}88, Swa76, Wil53a, Wil58c, Wil70, Wil75d, Wil88, Wil98c, Ano10, KT76, KT77, Wil49b, Wil90i]. Developments [Ano88, AWL^{+}88, Nee92, Wil75b, Wil80d]. Devices [Eng50]. diagnosis [Gil51a]. died [Ano10]. Difference [McL17, TW87]. Differential [Wil57d, Wil76d, Gil51b, LMW79, LJWB39, Wil40a, Wil56d]. Diffusion [AWL^{+}88, Fis50]. Digital [AWL^{+}88, Ben50, BW53a, Bow53, Bru20, CKW85, CK19a, CK19b, Cor58, Gil53, Ham58, Max57, McP52, Ran73, Ran75, Ran82, SW80, WWG51, WWG54, Wil56a, Wil56b, Wil57a, WWG57, Wil58a, Wil59a, Wil62c, Wil62d, Wil75a, Wil76e, WWG82, Wil86c, Wor53, BK52, CD55, CD57, Gil51b, Oet52, WS53, Wil53g, WRW58b, Wil60, WW79, WS82]. direct [CD55, CD57].

Discussion [UWE^{+}89, HNW^{+}48]. Discussions [BW53b, McP52]. Distributed [NH82, Wil78f, Wil78d, Wil79, WN80, Wil80e]. division [Wil56d]. Does [BHW^{+}90]. Domains [NW74]. Dominated [Wil64e, Wil65f]. d’opérations [Wil53e]. Doran [AWL^{+}88]. Dover [Wil64f]. Drafting [SW80]. dramatic [Wil83a, Wil91h]. Drawing [Wil62c]. Dream [Wil64f, Wil75g]. drive [Wil96d]. driven [Lan00, Wil96e]. Dutch [BWV^{+}88]. Dynamic [Wil65e, Wil00c]. dynamics [Wil73b].

E.D.S.A.C. [Wil53e, Wil89b, Wil89c]. Early [BWV^{+}88, CK80b, CK92a, CK19b, Cro90, Lav80b, Lav15, Nau80, Wil75b, Wil75c, Wil76e, Wil80d, Wil98c, Wil00b, Wil05, CK92c, KT76, KT77, Wil97d, Wil12, BWV^{+}88, WCK89]. Earth [Wil49c, Wil52b, Wil63b]. Eastern [Bro59]. Eckert [MHW87, Wil95d]. Ed [Wil72b, Wil74d, AMW^{+}88, AWL^{+}88, Wil95b]. edited [Wil64f]. Edition [Cor58]. Editorial [Wil91a]. Eds [Wil72b, AWL^{+}88, WLC86]. EDSAC [Wil86c, Ano15, Ano23, BBG^{+}50, BW53b, Bro53, CK80b, CK92c, CK19a, Dou79, Gil51a, Har99, Her24, Lav80a, Ley11, McL17, Nau51c, Ricxx, Smi13, Whe50, Whe51, Whe92b, Whe92a, Wil48, WR48, WR49, WR50a, WR50b, Wil50c, Wil52a, Wil53d, WPB54, WWG54, WW56a, WW56b, WW58, WR73, Wil75b, Wil76f, Wil77e, WR82, WWG82, WPB89, Wil90e, Wil92h, Wil94b, Wil97a, Wor50]. Education [Hop52]. EDVAC [SW80]. Edward [Wil87a, Wil97d]. effectuées [Wil53e]. efficiency [SW61]. Eigenvalue [Wil66a]. electric [WW59]. Electrical [WS86]. Electronic [AMW^{+}88, BS54, CKW85, Cop12, Cor58, Gil53, Ham58, Kid99, McP52, Nau51a, SW80, Wil49b, WR49, WR50b, WWG51, WWG54, WWG57, WS71, WR75, WWG82, Wil01b, Wil01c, Wil86c, BK52, Wil47a, WS53, WRW58b, WS82]. Elements [Wil67a, Wil67a]. Email [Wil90a]. Emergence [Cro92]. Emily [Wil64f]. EMPIRIC [Wil86b]. employs [Ano15]. Encyclopedia [BHK77, RM76]. end [Wil95c, Wil96b]. end-point [Wil95c, Wil96b]. Engine [Wil00e, Cop05, McL17, TW87]. engineer [Wil56b]. Engineering [Ben50, Tay52, Wil76g, Wil78e, WS86, Ari00]. engines [Wil91d, Wil91c, Wil64f]. England [Wil74d]. ENIAC [Ber00, MHW87, McL17, Wil13, Wil47a, WG82, Wil06]. Enigma [AWL^{+}88]. Environment [Wil72f]. EOL [Orl10]. Equations [Bro53, Wil57d, Gil51b, LMW79, Wil40a, Wil56d]. Errata [Ano88]. Essays [MHR80]. Establishment [GOB^{+}57, Pet70, Pet92]. establishments [VRW^{+}45]. etc [Orl10]. European [BWV^{+}88]. Evaluation [BS54, Tay52]. Evolution [CK19a, Wil98a]. exact [LMW79]. Examples [BN71, BSN82]. excitation [Wil51d]. execution [AEGP67]. Expanded [Cuf24]. Expectations [TW87, Wil91c]. Experience [WPB54, WPB89, Lan77]. Expériences [Wil53e, Wil53e]. experiment [Wil63a, Wil64h]. Experimental [BRW36]. Exploring [CK19b]. Extending [Les83].

F [Wil72b]. F.R.S. [Wil58b]. facilitate [Wil86b]. family [WU88]. famine [Wil90b]. Far [Wil98a]. Fast [Str59b]. Faster [Bow53]. Fastest [Cop12]. Fax [Wil90a]. February [BS96, HW04]. Fernando [Wil62e]. Field [Wil74b, Wil84c, Wil61c]. figures [Ano10]. Filing [Wil64g]. filled [Ano15]. fine [Wil86b]. First [AMW^{+}88, Kid99, RH00, Smi13, SW80, Ano15, Ash87, Lan77, Lan00, Wil96e, Wil00a, vdPWB82]. five [Wil90h]. ‘Flash [Wil13]. Floating [AEGP67, BW53b, Wil53g]. Floating-point [AEGP67]. Florida [IEE92]. Flowers [WG82]. Forbes [Kid99]. Forecasting [Wil70]. Foreword [Wil01e]. form [Wil83a, Wil91h]. formation [Wil40b]. Forschungsanstalt [VRW^{+}45]. Fortran [Wil93a]. Forum [Gro93]. forward [Wil65d]. Foundations [CK80a]. Fourier [BHW^{+}90, BK52]. Frame [Wil77a, Wil77b]. Frank [Wil61a]. Freiberger [Wil72b]. French [Wil53e, Wil67a]. FREng [HW04]. Frequencies [Fis50]. function [SW50]. Functions [BS54, WW57]. Further [BRW39]. Future [Lew21, Wil70, Wil74b, Wil77d, Wil77f, Wil78c, Wil84c, Wil88, Wil92g, Wil01d].

G [WG82, WS86]. Gabriel [Dwa18]. gamma [SW50]. gap [Wil01d]. Gardner [AWL^{+}88]. Gene [Fis50]. General [BS54, Wil50b, UWE^{+}89]. Genesis [Wil86a]. Genie [Kid99]. Geomagnetic [Wil62e]. Georgia [Wil78e]. German [AWL^{+}88, VRW^{+}45, dB05]. Gill [Cor58, CK11, Ham58, Wil86c]. Given [Dwa18, KT14, Mar17, Sco14, Ste16]. go [Smi13]. Gordon [Wil13]. grain [Wil86b]. Graphics [Wil67c]. grazing [Wil54]. Great [Lap96, Wil92d, Wil92f]. Greatest [Kid99]. Griffin [Wil61a]. Ground [Ber00]. Group [SW52]. Growth [Gor72, Wil69b]. Guest [Wil72b]. Guide [CDD^{+}63, WWG54].

H [WG82]. Half [Wil99a]. halt [Orl10]. Handbook [Wil62a]. handling [Nau51c]. Happenings [MHW87]. Hardware [Wil71f, Wil82a, Wil82b, Wil80f]. Harry [Wil62a]. Hartree [Ano88, Wil58b, AWL^{+}88, Wil84b]. Harvard [AWL^{+}88]. heights [SW52]. Held [GOB^{+}57, Wil97c, WK51]. help [Wil56b]. Helped [CRB24]. Herschel [Wil90i]. High [Eng50, Wil49e, Wil01a, CD55, CD57, Wil47a, Wil48, Wil49d, WW59, Wil01d]. High-Speed [Eng50, Wil49e, CD55, CD57, Wil47a, Wil48, Wil49d, WW59]. Hill [Wil62a]. Historical [Wil72c, Wil84a, Wil99b]. History [Anoxx, AWL^{+}88, BHW^{+}90, CK92a, Cer03, Cop06, Cuf24, Lav15, MHR80, RH00, Wil75c, Wil96d, dB05]. Howard [CWC99]. Huskey [Wil62a].

I. [HW03]. IBM [AEGP67, CRB24]. ICSE [Wil78e]. Idealised [Wor53]. Ideas [Lew21, UWE^{+}89]. IEE [Wil04b]. IEEE [Wil03, Asp92]. IFIP [Gil77]. II [CD57, Lip68]. Impact [Ano88, AWL^{+}88, Wil80e]. implementations [Wil82a, Wil82b]. important [Ano10]. Impressions [Nau80]. improving [SW61]. inaugural [WK51]. incidence [Wil54]. Incredible [Kid99]. Industry [Wil91a, Wil92b]. Influence [PHL92]. Influential [KT14]. Information [Cur12a, Gil77, Kid99, Weg68, Wil59b, Wil74d, Wil76b, Wil96d]. Infotech [Wil74d, Wil76b]. init.d [Orl10]. Initial [Ricxx, Whe50]. Inmos [AWL^{+}88]. inner [Wil68b]. Innovation [AWL^{+}88, O’R18a, Wil98a, Lan00, Wil96e]. Input [Wil57c]. Institution [VRW^{+}45]. instruction [Wil82c]. instructions [Wil83b]. Instruments [Ano88, AWL^{+}88, Wil84b]. integral [Wil54]. integrated [Wil13]. integration [Gil51b]. Integrity [Wil72d]. Intelligence [AMW^{+}88, Wil92e]. Interest [Gor72, Wil69b]. Interface [Wil71f, Wil00e]. International [KT14, Wil59b, Wil78e, GLL74]. Interview [And09]. Introduction [Bre68, CK92b, WWG54, Wil57c, Wil65b, Wil66b, Wil84b, Wil00e]. Introductory [Wil59c]. Invented [SW80]. Inventing [Ari00, Cer13]. Inventions [O’R18a]. Inverse [WW57]. Investigations [Wil75c, BRW36, BRW39]. ionic [SW52]. ionization [Wil39]. ionosphere [BRW36, BRW39, Wil36, Wil40b, Wil40c, Wil47b, Wil97d]. Ionospheric [Wil75c].

IRE [McP52]. ISBN [Wil95b]. ISCA [Wil03]. Issue [Lee92].

J [Cor58, Ham58, Wil74d, Wil86c]. J. [Ber84, WG82, WS86]. January [WK03]. jiden [Wil92c]. John [WU88, Wil01b, Wil01c]. Joint [Bro59, McP52, McP52]. Joseph [BHW^{+}90]. July [Wil03, WK51]. July-Aug [Wil03]. jump [Wil83b]. June [CK14, GOB^{+}57].

K. [WS86, WLC86]. kaisō [Wil92c]. KDF9 [OSWW65a, OSWW65b]. Keeping [Wil83b]. Kelly [WLC86, Wil95b, Wil96d]. Kemeny [WS86]. Key [IEE92]. Keynote [Wil82c]. Kilburn [WK03]. King [Wil87a]. Konpyūta [Wil92c]. Konrad [CRB24]. Kozaczuk [AWL^{+}88]. Kozyrakis [Ste16]. Kurtz [WS86].

L [Wil72b, WS86]. Laboratory [AWL^{+}88, Ano06, PHL92]. LAN [WH97]. Language [OSWW65a, OSWW65b, Wil98c, Wil63a, Wil64h, Wil68b, WM72, Wil86b]. Languages [Weg68, Wil93a, KT76, KT77, Wil64a]. Large [AWL^{+}88, Mil51, Str59b]. Large-Scale [AWL^{+}88]. last [Ano10, Hof11, Wil90h]. Later [Nee92]. Latitude [Wil63b]. learn [Oet52]. Lebenswerk [WLC86]. Lecture [WG82, Wil04b, Wil59c, Ryl58]. Lectures [CKW85, Ash87, WLC86]. LEO [Ari00, Cam03, Kid99, O’R18b, PHL92, Lan77]. letter [Gro93]. Letters [Wil03]. Library [BBG^{+}50, WWG82, Wil86c]. Life [BHW^{+}90, Wil95b]. Light [Wil93b]. like [Wil83b]. Lindgren [WS86]. line [Ano15, Wil64d]. Linear [BHW^{+}90, LMW79, Wil40a, Wil56d]. Lines [Kis92]. List [OSWW65a, OSWW65b, Wil63a, Wil64h]. list-processing [Wil63a, Wil64h]. Lists [Wil64c, Wil65c]. Literature [AMW^{+}88, AWL^{+}88, Wil65a, Wil69b]. Lives [Aus10]. Load [Wil71b, Wil71c]. Local [Les83, Wil80e]. Logical [Wil50b, Wil59c]. Loh [Dwa18]. London [Wil61a, Wil62a, Wil64f, Wil95b]. Long [Wil92g, BRW36, BRW39, Wil36, Wil40c, Wil47b, Wil56d]. long-division [Wil56d]. Long-Term [Wil92g]. Look [SG69]. Look-Ahead [SG69]. Ltd [Wil61a, Wil62a, Wil76b]. Ludgate [CRB24]. Luni [Wil62e]. Luni-solar [Wil62e]. Lyons [Wil01b, Wil01c].

M [AWL^{+}88, Gor72, Wil72b, WLC86, Wil95b]. M. [Bre68, Ham58, Max57, WLC86]. Machine [AWL^{+}88, Den99, Weg68, WR49, Wil49e, Wil53a, WWG54, Wil76a, Wil96a, BK52, Gil51b, Wil40a, Wil47a, Wil48, Wil49b, Wil49d, Wil51b, Wil81a, Wil89a, Wil21, Wil96d]. Machinery [BW53b, Wil56e, AWL^{+}88]. Machines [Ano88, AWL^{+}88, Bow53, GOB^{+}57, Wil51a, Wil51c, Wil53b, Wil53c, Wil57b, Wor53, dB05, HNW^{+}48, AWL^{+}88, Wil84b]. made [Wil13]. Magnetic [Wil57c, WW58, WW56a, WW56b]. magnetic-tape [WW56a, WW56b]. magneto [SW52]. magneto-ionic [SW52]. Maidenhead [Wil74d, Wil76b]. Main [SG69, Wil77a, Wil77b]. Mainframes [Wil78d]. Makin [CWC99]. Mallock [Wil40a]. Management [NW74, WH69, Wil84e, Wil74d]. Manchester [O’R18c, PL53, WK51]. Manual [AWL^{+}88]. March [HW04]. Marginal [WPB54, WPB89]. Mark [O’R18c]. Martin [Wil96d]. Masani [AWL^{+}88]. Maschinen [dB05]. masiny [Wil60]. Mass [Cor58]. master [Cop05]. Mathematical [Ano88, AWL^{+}88, Kid99, PHL92, Wil90c, vdPWB82, Wil50c, Ano06]. Mathematics [Gil53, Ber84]. Matrices [Wil67a]. Maturing [Wil91a]. Mauchly [SW80]. Maurice [Ano10, Aus10, CK14, Cor58, Dwa18, Hen86, KT14, Mar17, Sco14, Ste16, Wil13, Wil86c, And09, Ano10, Ben12, Cur12b, Gil83, Gro93, Hof11, Lee93, Lee95]. Maverick [AWL^{+}88]. Maxwell [Ber84, Wil84d]. May [Wil78e, Wil97c, Wil56b]. McGraw [Wil62a]. McGraw-Hill [Wil62a]. McLean [AWL^{+}88]. Mechanical [Kis92]. Meeting [vdPWB82]. Meetings [WG82]. Mein [WLC86]. Memoirs [AMW^{+}88, Wil85, Hen86]. Memories [Cuf24, Dou79, Wil65e, Wil71h, Wil00c, Wil01d]. Memory [SG69, Wil72f, Wil76h, Ano15, Ley11, WR48, Wil82a, Wil82b, Wil95c, Wil01a, Wil01d]. Mercury [Ano15, Ley11]. mercury-based [Ley11]. Mercury-filled [Ano15]. Method [WW57, GHR57, Wil40a]. Methods [Wil61a]. Michael [HW04]. Micro [WS53, Wil03, Wil77g]. Micro-programming [WS53, Wil77g]. Microchip [AMW^{+}88]. Microcomputers [Wil77a]. Microelectronics [AWL^{+}88]. microprocessors [Wil77b]. Microprogram [Tuc67]. Microprogrammable [Wil78b]. Microprogramming [Wil59d, Wil69b, WS71, Wil71g, Wil75d, WS82, Wil86a, BHW^{+}90, Gor72]. Milestones [Bru20]. Mind [AWL^{+}88]. Minicomputers [Wil78d]. Minimalist [GL03]. minor [Nau51b]. mistakes [Gil51a]. Model [AEGP67, Lip68, Wil69c, WN80]. Modern [Cer03, Cop06, Cop05]. Module [SG69]. molecular [Ken63]. month [Ano10]. Moore [MHW87, WLC86, CKW85]. Morrison [Wil64f]. most [Ano10]. motion [Nau51b]. Mr [Wil95b]. Mr. [Wil83a, Wil91h]. MSRC [Ber00]. Multi [SG69, Wil64b]. multi-access [Wil64b]. Multi-Module [SG69]. Multiple [Wil67b, Wil68a, Wil73a, Wil67e, WN68]. Multiple-Access [Wil67b, Wil73a, Wil67e, WN68].

Multiprogramming [Wil72f]. Museum [WLC86, MHW87]. Myoglobin [Ken63].

N [WS86]. name [Nau51c]. nature [Ryl58]. Needham [HW04, HJ04]. Nets [Wil02a]. Network [Les83]. Networks [Wil90a]. Neumann [SW80]. Neural [Wil02a]. No [Wil72b, Wil74d, Wil76b, Wil92c]. Norbert [AWL^{+}88]. Norris [AWL^{+}88]. Note [Wil56c, WW57]. Notebooks [Wil81b]. Notes [WW57, BHW^{+}90, vdPWB82]. November [Ber00, CK14]. Numbers [CWC99, Mil51]. Numerical [Bre68, Wil57d, Wil65b, Wil66b, Wil67a]. Numérique [Wil67a].

Obituary [HW03, Wil58b]. Object [Wil93a]. Object-Oriented [Wil93a]. oblique [Wil47b]. Office [VRW^{+}45, Wil62c, Wil01b, Wil01c]. omit [Ley11]. On-line [Wil64d]. one [Ano10]. Online [Wil65d]. Operating [IEE92, WN79, Wil92g, Wil94a, Wil92a]. Operation [AWL^{+}88, SG69, Wil64d, PL53]. Operational [Wil53e]. Operations [BW53b]. Optimum [Pet70, Pet92]. order [Wil40a]. Orders [Ricxx, Whe50, Wil53g]. Ordinary [Wil57d]. Organization [Weg68, Whe50]. Organizational [AWL^{+}88]. Oriented [Wil93a]. Origins [Ran73, Ran75, Ran82, Wil94b]. Oscillations [Wil49a, Wil49c, Wil63b, WW47, Wil51d, Wil52b]. Other [AMW^{+}88, AWL^{+}88]. our [Cer13]. outer [Wil68b]. Outlook [Wil67d, Wil61c]. Output [Wil57c]. Overseas [Wil99c]. Overview [Wil84a].

Pa [McP52, Bro59]. Padé [LMW79]. pages [Wil62a, Wil72b, Wil74d, Wil76b]. paging [Wil73b]. paionia [Wil92c]. Panel [Wil03]. Paper [Ano08, CRB24, KT14]. Papers [Lap96, Lew21, McP52, Ran73, Ran75, Ran82, Swa76, AWL^{+}88, WW57]. parabolic [LMW79]. Park [Ley11, Smi13]. Part [Lip68, Wil96c, WN68]. Partha [Sco14]. Parthasarathy [Sco14]. partial [LMW79]. parts [Smi13]. Passages [Wil95b, Wil95b]. Past [Cer13, Wil74b, Wil84c, Wil88]. Patent [CRB24]. Peacock [Wil90i]. Pencil [Bly57]. Pendulum [Wil98b]. Pennings [AWL^{+}88]. Pennsylvania [BS96]. People [Lav80b]. Percy [CRB24]. performance [WR47, Wil01a, Wil01d]. Personal [Cam03, Wil84e, Wil88, Wil02a, Har99, Lan77, Wil92a]. Perspective [Asp92, GL03, Wil90h]. Perspectives [Wil72c, Wil90f, Wil91e, Wil91f, Wil92e, Wil92f, Wil92g, Wil93a, Wil93b, Wil95a, Wil90g, Wil91g, AMW^{+}88]. Perturbations [Nau51a]. Philadelphia [BS96, Bro59, McP52]. Philip [Wil64f]. Philosopher [Wil95b, Wil95b]. Philosophical [AMW^{+}88]. photo [WW59]. photo-electric [WW59]. Physics [Ano88, AWL^{+}88, Gil53]. Pickering [Wil95b]. Pinkerton [Wil01b, Wil01c]. Pioneer [Wil71d, Wil85, Wil91b, Hof11, Orl10, Wil77c, AMW^{+}88, BHW^{+}90, Hen86]. pioneering [Ley11]. Pioneers [Gla97]. pipeline [Wil83b]. Pivotal [O’R18a]. Place [Wil62d]. planet [Nau51b]. Plans [Wil00e, UWE^{+}89]. point [AEGP67, Wil95c, Wil96b]. Portrait [AWL^{+}88]. Position [SW80]. possible [Wil13]. Post [VRW^{+}45]. Potential [Wil71g]. power [Wil83c]. Pp [Wil95b, Wil96d, Cor58, Gor72, Wil03]. Practical [Wil61d, Ano10, Wil13, Wil48]. Practice [CK19a]. praise [CK11]. Pray [Wil83a, Wil91h]. Predicted [Wil91e]. Preparation [BBG^{+}50, Cor58, Ham58, WWG51, WWG54, WWG57, WWG82, Wil86c]. Present [Wil71g, Wil74b, Wil84c]. Presented [McP52]. Preserving [Wil72d]. Presper [Wil95d]. Press [Wil72b]. Prestige [Wil04b]. Price [Wil72b, Wil74d, Wil76b]. Prime [Mil51]. Principal [Swa76]. Principles [BSN82, Wil75d]. Prize [Ano81]. problem [WH97, Wil66a]. Problems [Gil53]. Proceedings [BS96, Bro59, GOB^{+}57, GLL74, Wil78e, AWL^{+}88, Gil77, IEE92]. process [Gil51b]. Processes [NW74]. Processing [Gil77, GOB^{+}57, OSWW65a, OSWW65b, Wil59b, Wil78f, Wil79, Ken63, Wil63a, Wil64h]. Processor [SG69, Wil97b, Wil82c, Wil90g]. production [Smi13]. Prof. [Wil58b]. Professor [Ano10, Aus10]. Program [Ricxx, WR75, Ano15]. Programme [Str59a, Whe50, Wil49d]. Programme-controlled [Str59a]. Programmer [CDD^{+}63, Wil64g, Wil00e, Wil91g]. Programmes [BBG^{+}50, Gil51a]. Programming [CK80a, CK80b, CK92c, Nau80, Weg68, Whe92a, WWG54, Wil74a, Wil76g, Wil80d, Wor53, KT76, KT77, WS53, Wil53f, Wil64a, Wil68b, Wil77g, Wil86b, BHW^{+}90, Oet52]. Programs [Cor58, Ham58, WWG51, WWG54, WWG57, WWG82, Wil86c]. Progress [Wil49e, Wil92b, Wil04b, Wil90h]. Project [PHL92]. projects [UWE^{+}89]. Prometheus [Kid99]. proposals [SW61]. Prospective [Ano88, AWL^{+}88]. Protection [NW74, Wil84e, Wil82a, Wil82b, Wil86b]. Proteins [Ken63]. provides [Ano15]. Proving [Ber00]. Publications [Ham58, Max57]. Publishing [Cor58, Wil62a]. Pure [Wil53f]. put [Wil99b].

Quadrature [Wil56c]. Queisser

[AMW^{+}88]. Queries [BWV^{+}88, BHW^{+}90, BW00b, Kis92, SW80, Wil00d]. quest [Wil00a]. Question [AMW^{+}88].

R [Wil58b, Wil84b, WLC86]. radar [WR47].

Radio [Bly57, Ryl58]. Rajwar [KT14]. Ranganathan [Sco14]. Rapport [Wil50b]. Ravi [KT14]. Razdeleniem [Wil72e, Uil72]. reaches [Orl10]. Read [AWL^{+}88]. readers [WW59]. Reading [Cor58, Wil81b]. Readings [BN71, HJS00]. Real [Wor53]. rebuild [Ley11, Smi13]. Rechentechnik [dB05]. Rechnen [dB05]. reciprocal [SW50]. Recollections [Ano23, Gla97, Wil87a, Wil00b, Wil05, Wil12]. Reconstructing [Her24]. Record [Ber00]. Rectangle [Wil67a]. Rectangles [Wil67a]. Reference [WWG82, Wil86c]. reflected [BRW36, BRW39]. reflection [Wil36]. Reflections [Wil81b, Wil84d]. reflexion [Wil40c, Wil47b]. Region [Wil40b, Wil39]. Reichspost [VRW^{+}45]. related [Wil96b]. relations [GHR57]. Remarks [Wil57d]. Remember [Wil06]. Remembered [Wil97c, Aus10]. Reminiscence [Cam03]. reminiscences [Har99]. Repairing [Wil61e, Wil61d]. replica [Smi13]. Report [AWL^{+}88, Wil76b, WG82, Wil74d, SW80]. Reports [VRW^{+}45]. reprint [Wil84b]. Rescuing [Kid99]. Research [Cro92, GOB^{+}57, VRW^{+}45, Wil92b, Wil97d]. resonance [WW47]. Response [BHW^{+}90]. Retrospect [Wil73a, WG82]. reveal [Ken63]. Review [Cor58, Gor72, Hen86, McP52, Wil61a, Wil62a, Wil66a, Wil67a, Wil69a, Wil72b, Wil74d, Wil76b, Wil78a, Wil78b, Wil95b, Wil00a, Wil86c, Wil96d]. Reviews [AMW^{+}88, AWL^{+}88, Bre68, Kid99, Wil64f, WS86, WLC86, Ano88]. Revised [Cor58, Cuf24]. Revisionist [Wil98c]. Revisitation [Wil02a]. Revisiting [Wil91f]. Revolution [AWL^{+}88]. Ring [Wil75a, WW79, BS82, Nee92]. RISC [Wil83b]. RISC-like [Wil83b]. rise [Wil90g]. Roger [HJ04, HW04]. Role [SW80]. routines [BBG^{+}50]. routing [WPB89]. Routining [WPB54]. Rowland [AWL^{+}88]. Rubinoff [Wil72b]. Russian [Uil72, Wil60, Wil72e].

s [Uil72, Wil72e]. SA [GOB^{+}57]. Saga [AWL^{+}88]. Salisbury [GOB^{+}57]. Sampling [Wil61a]. San [Wil62e]. Scale [AWL^{+}88]. Scanning [Cer13]. Scene [Wil71e]. Schonland [Aus01]. School [WLC86, Wil87a, CKW85, MHW87]. Science [AWL^{+}88, BS96, BHK77, BHW^{+}90, Cro92, Lap96, Lew21, RM76, WS86, Ber84, Wil99b]. Scientific [Cro90]. Scientist [Aus01]. Searching [Wil74a]. Second [Wil58c, Wil40a]. Security [Wil84e, Wil90d, Wil91f]. Segmentation [Wil71h]. Select [O’R18a]. Selected [Ran73, Ran75, Ran82]. Selection [Fis50]. Self [OSWW65a, OSWW65b, Wil61d, Wil61e, Wil63a, Wil64h]. Self-Compiling [OSWW65a, OSWW65b, Wil63a, Wil64h]. Self-Repairing [Wil61e, Wil61d]. September [Gor72, GLL74]. Sequence [AWL^{+}88]. Serial [Wor53]. Servers [Wil98a]. sesquicentennial [Ber84]. Session [Wil57c]. set [Wil82c, WLC86]. SHARE [vdPWB82]. shared [Wil64d]. Sharing [CDD^{+}63, Str59b, Wil69c, Wil71b, Wil71c, Wil75e, Orl10, Str59a, Uil72, Wil65d, Wil68c, Wil68d, Wil72g, Wil75f, Wil72e]. Sherwood [Mar17]. ship [WR47]. Short [Bre68, WW57, Wil66b]. sic [Wil56c]. sign [GHR57]. silicon [Wil90e]. simple [Wil63a, Wil64h]. Simplified [SG69]. Simultaneous [SG69, Wil40a]. Sir [Ano10, Aus10, CK14, Wil13, Wil97d]. Sistemy [Uil72, Wil72e]. Size [Pet70, Pet92, Wil83c]. sketch [Wil86b]. Slave [Wil65e, Wil71h, Wil00c]. Slide [WLC86]. small [LJWB39]. Society [Kid99]. Sociological [AMW^{+}88]. Software [Gla97, WH69, Wil71f, Wil76g, Wil78e, Wil90f, Wil91g]. Solar [Wil62e]. Soldier [Aus01]. Solitons [Wil93b]. Solution [Bro53, Wil56d, Wil57d, LMW79, WH97]. solving [Wil40a]. Some [Ben50, Dou79, SW61, Wil57d, Wil74c]. Sorting [Wil74a]. sources [Ryl58]. Space [Wil69c]. SPEC [Wil03]. Special [Nau51a, WWG82, Wil86c, Wil59c]. Specialized [Wil04a]. Species [WH69]. Speed [Eng50, Wil49e, CD55, CD57, Wil47a, Wil48, Wil49d, WW59, Wil83c]. Square [Wil67a]. Squares [Ricxx]. St [WU88]. Stand [BHW^{+}90]. Stanley [Cor58, Ham58, Wil86c]. start [Smi13]. State [Wil76b, Wil74d]. statements [Wil64a]. stations [Wil92a]. Status [Wil71g]. step [Gil51b, Wil65d]. step-by-step [Gil51b]. Storage [WR50b, Wil57c, WW58, Wil65e, Wil00c, Ano15, WW56a, WW56b]. Stored [WR75, Ano15]. stored-program [Ano15]. Stories [Kid99]. Story [BHW^{+}90, Kid99, Lav80b, AMW^{+}88]. Stourbridge [Wil87a]. Strip [Wil56c]. Structural [Lip68]. Structure [Ken63]. Structured [Wil76g]. Structures [BN71, BSN82, Weg68, CD55, CD57, Wil71a, Wil72a]. Struggle [Cop12, Cop05]. studies [Wil99b]. Study [BHW^{+}90, Lan00, Wil83a, Wil91h]. Sub [BBG^{+}50]. Sub-routines [BBG^{+}50]. subroutine [Nau51c]. Subroutines [WWG82, Wil86c]. successor [Ano06]. Supplement [WWG54]. Supply [Kid99]. support [Wil82a, Wil82b]. Surprises [Wil99a]. Survey [Wil69b]. Surveys [Gor72, Wil61a]. Swing [Wil98b]. Switzerland [GLL74]. Symposium [AWL^{+}88, Ber00, Bow53, GLL74, KT14, Ber84, Wil97c]. syntax [Wil68b]. synthesis [BK52]. System [AEGP67, NH82, Wil64g, Wil69c, WH69, WN79, Wil53g, WW56a, WW56b, Wil64b, Lip68, Tuc67, CDD^{+}63, Wil73a, WN80]. System/360 [AEGP67, Tuc67, Lip68]. systematic [GHR57]. Systems [HJ04, IEE92, Whe92a, Wil61d, Wil68a, Wil68c, Wil68d, Wil71b, Wil71c, Wil72e, Wil72g, Wil75f, Wil77f, Wil78d, Wil92g, Ari00, Uil72, Wil67e, WN68, Wil80e, Wil92a, Wil94a, Wil01a, Wil76b].

T [WG82, WS86]. Table [SW50, Wil54]. Tables [Wil90c]. tabular [Wil71a, Wil72a]. Take [Wil98a]. Talking [Den99]. Tape [CK92a, Wil57c, WW58, WW56a, WW56b, WW59]. targets [WR47]. Teaching [Cro92]. tech [Wil13]. Technical [WW57]. Techniques [CKW85, Wil61b, Wil62b, Ken63]. Technology [AWL^{+}88, BHK77, HJ04, Kid99, Wil96d, dB05, Ber84, Wil99b]. Temperature [Wil63b]. Ten [Wil77g]. Term [Wil92g]. Their [Ano88, AWL^{+}88]. Them [Lav80b]. Theoretical [Wil39]. Théorie [Wil67a]. Theory [CKW85, HJ04, Kid99, Wil67a, SW52, WW47, Wil40b, Wil40c, WR47]. thermal [Wil51d]. Think [Wil51c, Wil53b, Wil53c, Wil57b]. Third [IEE92]. Thought [Bow53]. Thoughts [Wil74c]. Thwart [CRB24]. Time [CDD^{+}63, Str59b, Wil64d, Wil68c, Wil68d, Wil69c, Wil71b, Wil71c, Wil72g, Wil75e, Wil75f, Ano23, Orl10, Str59a, Wil13, Wil65d, Uil72, Wil72e]. Time-shared [Wil64d]. Time-Sharing [CDD^{+}63, Wil69c, Wil71b, Wil71c, Wil68c, Wil68d, Wil72g, Wil75f, Wil72e, Uil72]. Timeline [Anoxx]. Timothy [Mar17]. TITAN [Ano23, Nee92]. Today [Wil77h]. Tom [WK03]. Tomorrow [Wil61b, Wil62b]. topics [Wil96b]. Toronto [Gil77]. trailblazer [Smi13]. Translation [Wil56e]. transmission [Wil61c]. tribute [HJ04, Wil95d]. true [Wil75g]. Tube [Wil00b, Wil05, Wil12]. Turing [Ash87, AWL^{+}88, Ben12, Cop05, Cop12]. Twentieth [MHR80]. twenty [Ash87]. Two [AWL^{+}88, BWV^{+}88]. Type [Bly57, Wil64a].

Ubiquitous [Wil95b]. UK [Ano10]. Ultrasonic [Wil76h, WR48]. uncle [Wil92d, Wil92f]. Unit [Wil81b, AEGP67, WR48, WRW58a, WRW58b]. University [CK19a, PL53, WK51, CK80b, CK92b, CK92c]. unknown [Bro49, Str54]. USA [BS96, Wil78e]. Use [BBG^{+}50, Wil56c, Wil72f, WWG82, Wil03, Wil86c, CD55, CD57, Wil50c, Wil53g]. Useful [Wil64c, Wil65c]. User [WWG54, Wil96e, Lan00]. User-driven [Wil96e, Lan00]. uses [Wil99b]. Using [Wil75a, LMW79, Nau51c, Wil40a]. Utility [Wil64g].

V [Ano10, Bre68, Cor58, Gor72, Ham58, Lee93, Max57, Wil86c, VRW^{+}45]. V1 [Gor72]. Vacuum [Wil00b, Wil05, Wil12]. Variation [Wil62e, Wil63b]. Vector [BS54, Nau51c]. Vector-Coupling [BS54]. vector-handling [Nau51c]. velocities [SW52]. very [BRW36, BRW39, Wil36, Wil40c, Wil47b, Wil65d]. View [Wil99c]. Vincent [CK14, Lee95]. Vintage [Lav80b]. vistas [Ber84]. Visual [dB05]. VLSI [Wil90g, Wil91a]. Vol [Wil72b]. Volume [Ano08, Wil74a]. Vremeni [Wil72e, Uil72]. vycislitel’nye [Wil60].

W [Wil72b]. wall [Wil95c]. War [AWL^{+}88]. Was [AWL^{+}88, BW00b, BW00a, Wil00d, Ano10]. Wave [BS54]. waves [BRW36, BRW39, Wil36, Wil40c, Wil47b]. Way [Wil76a, Wil96a, Wil51b, Wil81a, Wil89a, Wil21]. Weapons [GOB^{+}57]. Wesley [Cor58]. Wheeler [CK11, Cor58, Ham58, Wil86c]. which [Wil99b]. Whiggism [BHW^{+}90]. Whirlwind [BHW^{+}90, BHW^{+}90, Eve52, Tay52]. Who [Lav80b, SW80, Ano10]. wide [Wil52b, Wil80e]. wide-band [Wil80e]. Width [Wil56c]. Wiener [AWL^{+}88].

Wildes [WS86]. Wilkes

[CK14, Cor58, Dwa18, Gor72, Ham58, KT14, Mar17, Sco14, Ste16, Wil86c, CK11, And09, Ano81, Ano85, Ano90, Ano91, Ano08, Ano10, AMW^{+}88, Aus10, Ben12, Bre68, Cur12b, Gil83, Gro93, HT81, Hen86, Hof11, Lee93, Lee95, Max57, Orl10, Wil13]. Will [Ley11]. William [Wil96d]. Williams [WLC86]. Winthrop [WU88]. wireless [BRW36, Wil36, Wil40c, Wil47b]. Wirukusu [Wil92c]. Wishes [Lee93]. WISP [OSWW65a, OSWW65b]. Work [BHW^{+}90, Wil92a]. Workers [Kid99]. Works [AWL^{+}88]. Workshop [IEE92]. Workstation [IEE92]. Workstations [Wil98a]. World [AWL^{+}88, Cop12, Kid99, Wil52b, Wil64e, Wil65f, Wil90d, Wil91f, Wil94a, Wil96e, Wil02b]. World-wide [Wil52b]. Worthy [AWL^{+}88]. Writable [Wil72f]. written [Nau51c].

XII [BS54].

Yates [Wil61a]. Year [Wil92e]. Years [Ber00, Ash87, Wil77g, Wil90h]. York [Wil64f, Wil72b].

Zuse [WLC86, CRB24].

References

Anderson:1967:ISMb

[AEGP67] S. F. Anderson, J. G. Earle, R. E. Goldschmidt, and D. M. Powers. The IBM System/ 360 Model 91: Floating-point execution unit. IBM Journal of Research and Development, 11(1):34–53, January 1967. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5392016.

Aspray:1988:RBB

[AMW^{+}88] William Aspray, A. R. Mackintosh, Eric A. Weiss, John McCarthy, and M. Woodger. Reviews: Burks and Burks: The First Electronic Computer: The Atanasoff Story; Queisser: The Conquest of the Microchip; Bloomfield (ed.): The Question of Artificial Intelligence: Philosophical and Sociological Perspectives; Wilkes: Memoirs of a Computer Pioneer; other literature. Annals of the History of Computing, 10(3):221–233, July/ September 1988. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/
an/books/an1988/pdf/a3221.
pdf;   http://www.computer.
org/annals/an1988/a3221abs.
htm.

Anderson:2009:IIM

[And09] David P. Anderson. Interview: An interview with Maurice Wilkes. Communications of the ACM, 52(9):39–42, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1981:WP

[Ano81]  Anonymous.  The Wilkes

Prize. The Computer Journal, 24(1):1, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        01/tiff/1.tif.

Anonymous:1985:WA

[Ano85]  Anonymous.  The Wilkes

Award 1984. The Computer Journal, 28(5):507, 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.org/
                                        content/28/5/507.full.pdf+
                                        html.

Anonymous:1988:ERH

[Ano88]  Anonymous.  Errata: Re-

views: Hartree: Calculating Machines: Recent and Prospective Developments and Their Impact on Mathematical Physics and Calculating Instruments and Machines, 10(1)

93. Annals of the History of Computing, 10(3):234, July/ September 1988. CODEN AH-COE5. ISSN 0164-1239. URL

                                        http://dlib.computer.org/
                                        an/books/an1988/pdf/a3234.
                                        pdf;   http://www.computer.
                                        org/annals/an1988/a3234abs.

htm. See [AWL^{+}88].

Anonymous:1990:WA

[Ano90] Anonymous. Wilkes Award. The Computer Journal, 33(2): 163, 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.org/
                                        content/33/2/163.1.full.pdf+
                                        html.

Anonymous:1991:WA

[Ano91] Anonymous. Wilkes Award.

The Computer Journal, 34(2): 112, 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/34/2/112.full.pdf+
html.

Anonymous:2006:AML

[Ano06] Anonymous. Archives of the Mathematical Laboratory and its successor, the Computer Laboratory, 1934–2006. Cambridge University Library, 2006.

URL https://archivesearch.
lib.cam.ac.uk/repositories/
2/archival_objects/7939. ef-

erence Code: GBR/0265/UA/COMP.

Anonymous:2008:WBP

[Ano08] Anonymous. The Wilkes Best Paper Award in volume 49 (2006). The Computer Journal, 51(1):??, January 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
cgi/reprint/51/1/NP.

Anonymous:2010:PSM

[Ano10]  Anonymous.  Professor Sir

Maurice Wilkes (1913–2010): Maurice V. Wilkes, who died last month aged 97, was one of the most important figures in the development of practical computing in the UK. Web site., December

2010. URL https://www.cam.

                                        ac.uk/news/professor-sir-
                                        maurice-wilkes-1913-2010.

Anonymous:2015:ECE

[Ano15]  Anonymous.  1949: ED-

SAC computer employs delayline storage: Mercury-filled delay line provides memory for first stored-program computer. Web site, August 24, 2015. URL https:/

                                        /www.computerhistory.org/
                                        storageengine/edsac-computer-
                                        employs-delay-line-storage/

.

Anonymous:2023:RTE

[Ano23] Anonymous. Recollections of time with EDSAC & TITAN. Web site., 2023. URL https://

                                        retrocomputingforum.com/t/
                                        recollections-of-time-with-
                                        edsac-titan/4401.

Anonymous:20xx:TCH

[Anoxx] Anonymous. Timeline of computer history. Web site.,

                                        20xx.  URL https://www.
                                        computerhistory.org/timeline/
                                        computers/.

Aris:2000:ISE

[Ari00] J. Aris. Inventing systems engineering [LEO]. IEEE Annals of the History of Computing, 22(3):4–15, July/September 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        85/18655/00859522.pdf.

Ashenhurst:1987:ATA

[Ash87] Robert L. Ashenhurst, editor.

ACM Turing Award Lectures: the first twenty years, 1966– 1985. ACM Press anthology series. ACM Press and Addison-Wesley, New York, NY, USA and Reading, MA, USA, 1987. ISBN 0-201-07794-9. xviii + 483 pp. LCCN QA76.24 .A33 1987.

Aspray:1992:IPc

[Asp92] William Aspray. From IEEE’s perspective. IEEE Annals of the History of Computing, 14(4): 5–??, October/December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://dlib.
computer.org/an/books/an1992/
pdf/a4005.pdf.

Austin:2001:SSS

[Aus01] Brian Austin. Schonland: Scientist and Soldier. Institute of Physics Publishing, Bristol, UK and Philadelphia, PA, USA, 2001. ISBN 0-7503-0501-0 (hardcover), 1-4200-3357-3 (PDF e-book). xviii + 639 pp. LCCN QC16.S2646 A97 2001; Q143.S36 A87 2001. URL http:

//www.crcnetbase.com/isbn/
9781420033571.

Austin:2010:LRP

[Aus10] Brian Austin. Lives remembered: Professor Sir Maurice Wilkes. The Times [London, UK], ??(??):??, December 9, 2010. ISSN 0140-0460, 0956-1382. URL

https://www.thetimes.com/
article/lives-remembered-
professor-sir-maurice-wilkes-
rx3x5mcvm9n?msockid=11b880866543636c27ec94db64796211.

Aspray:1988:RCD

[AWL^{+}88] William Aspray, Maurice V. Wilkes, Albert C. Lewis, Greg Mellen, Harold Chucker, Robert V. D. Campbell, Wendy Wilkins, G. J. Tee, Ernest Braun, and Arthur L. Norberg. Reviews: Carpenter and Doran (eds.): A. M. Turing’s ACE Report of 1946 and Other Papers; Masani (ed.): Norbert Wiener: Collected Works with Commentaries; Kozaczuk: Enigma: How the German Machine Cipher Was Broken and How It Was Read by the Allies in World War Two; Worthy: William C. Norris: Portrait of a Maverick; Harvard Computation Laboratory: A Manual of Operation for the Automatic Sequence Controlled Calculator; Proceedings of a Symposium on Large-Scale Digital Calculating Machinery; Gardner: The Mind’s New Science: A History of the Cognitive Revolution; Hartree: Calculating Machines: Recent and Prospective Developments and Their Impact on Mathematical Physics and Calculating Instruments and Machines; McLean and Rowland: The Inmos Saga; Pennings and Buifendam (eds.): New Technology as Organizational Innovation: The Development and Diffusion of Microelectronics; other literature. Annals of the History of Computing, 10(1):80–97, January/ March 1988. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/
an/books/an1988/pdf/a1080.
pdf;   http://www.computer.
org/annals/an1988/a1080abs.

htm. See minor erratum [Ano88]: Hartree as a mathematical physicist, not a physical chemist.

Bennett:1950:PPE

[BBG^{+}50] J. M. Bennett, R. A. Brooker, S. Gill, D. R. Hartree, E. N. Mutch, B. Noble, J. P. Stanley, D. J. Wheeler, M. V. Wilkes, and B. H. Worsley. Preparation of programmes for the ED-SAC and the use of the library of sub-routines. Report, University Mathematical Laboratory, Cambridge, UK, September 1950. 41 + 26 + 39 + 11 pp.

Bennett:1950:SEA

[Ben50] John Makepeace Bennett. Some Engineering Applications of Digital Computation. Ph.D. thesis, University of Cambridge, Cambridge, UK, 1950. URL

https://www.mathgenealogy.
org/id.php?id=110514.

Bentley:2012:MWA

[Ben12]  Peter J. Bentley.  Mau-

rice Wilkes on Alan Turing. OUP Blog, June 2012. URL

http://blog.oup.com/2012/
06/maurice-wilkes-on-alan-
turing/.

Berger:1984:JCM

[Ber84] Melvyn S. (Melvyn Stuart) Berger, editor. J. C. Maxwell, the sesquicentennial symposium: new vistas in mathematics, science, and technology. North-Holland, Amsterdam, The Netherlands, 1984. ISBN 0-444-86707-4 (Elsevier). LCCN QC669 .J17 1984.

Bergin:2000:YAC

[Ber00] Thomas J. Bergin, editor. 50 Years of Army Computing: From ENIAC to MSRC: A Record of a Symposium and Celebration November 13 and 14, 1996 Aberdeen Proving Ground, volume ARL-SR-93. U.S. Army Research Laboratory, Aberdeen, MD, USA, September 2000. ISBN 0-9702316-1-X. LCCN QA76.5 .S95 1996. URL https:/

                                        /apps.dtic.mil/sti/tr/pdf/
                                        ADA431730.pdf;   https://
                                        permanent.fdlp.gov/lps58495/
                                        lps58495.pdf.

Belzer:1977:ECS

[BHK77] Jack Belzer, Albert G. Holzman, and Allen Kent, editors. Encyclopedia of Computer Science and Technology. Marcel Dekker, New York, NY, USA, 1977. ISBN 0-8247-2251-5. various pp. LCCN QA76.15 .E5.

Buchholz:1990:CQDa

[BHW^{+}90] Werner Buchholz, R. Anthony Hyman, Maurice V. Wilkes, Richard E. Smith, George T. Jacobi, Alice R. Burks, Arthur W.

Burks, I. Bernard Cohen, I. O. Grattan-Guinness, S. I. Gass, and Mike Woodger. Comments, queries, and debate: Whiggism in the History of Science and the Study of the Life and Work of Charles Babbage; Whirlwind and Microprogramming; More on Wilkes, Whirlwind, and Microprogramming; Pioneer Biographies; The Atanasoff Story — A Response; Notes on Babbage, Aiken, and Bowditch; Joseph Fourier’s Anticipation of Linear Programming; What Does BNF Stand for? Annals of the History of Computing, 12(1):62–72, January/ March 1990. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/
an/books/an1990/pdf/a1062.
pdf;   http://www.computer.
org/annals/an1990/a1062abs.
htm.

Bennett:1952:CFS

[BK52] J. M. Bennett and J. C.

Kendrew. Computation of Fourier synthesis with a digital electronic calculating machine. Acta Crystallographica, 5(Part 1):109–116, January 1952. CODEN ACCRA9. ISSN 0365-110X (print), 1600-8642 (electronic). URL https:

//journals.iucr.org/paper?
S0365110X52000228.

Blythe:1957:NTP

[Bly57] J. H. Blythe. A new type of pencil beam aerial for radio astronomy. Monthly Notices of the Royal Astronomical Society, 117(6):644–651, December 1957. CODEN MNRAA4. ISSN 0035-8711 (print), 1365-2966 (electronic).

Bell:1971:CSR

[BN71] C. Gordon Bell and Allen Newell, editors. Computer Structures: Readings and Examples. McGraw-Hill, New York, NY, USA, 1971. ISBN 0-07-004357-4. xix + 668 pp. LCCN TK7888.3 .B4.

Bowden:1953:FTT

[Bow53] Baron Bertram Vivian Bowden, editor. Faster Than Thought: a Symposium on Digital Computing Machines. Sir Isaac Pitman and Sons, Ltd., London, UK, 1953. LCCN QA76.5 .B66.

                                        URL https://archive.org/
                                        details/FasterThanThought.

With a foreword by the Right Honourable the Earl of Halsbury.

Brenner:1968:RSI

[Bre68] Joel Brenner. Reviews: A Short Introduction to Numerical Analysis, by M. V. Wilkes. American Mathematical Monthly, 75 (9):1031, November 1968. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Brooker:1949:U

[Bro49] Ralph Anthony (Tony) Brooker. [unknown]. Ph.D. thesis, University of Cambridge, Cambridge, UK, 1949. URL https:/

                                        /www.mathgenealogy.org/id.
                                        php?id=297145.

Brooker:1953:SAE

[Bro53] R. A. Brooker. The solution of algebraic equations on EDSAC. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 48(2):255–270, April 1953. CODEN PCPSA4. ISSN 0008-1981. URL https:

//www.cambridge.org/core/
journals/mathematical-proceedings-
of-the-cambridge-philosophical-
society/article/solution-
of-algebraic-equations-on-
the-edsac/01E6A530985EBAE6848CCB5958258649.

Broomall:1959:PEJ

[Bro59] John M. Broomall, editor. Proceedings of the Eastern Joint Computer Conference, December 3–5 1958, Philadelphia, PA, volume 18 of AIEE Conference Proceedings. American Institute of Electrical Engineers, New York, NY, USA, July 1959. ISBN 1-4503-7866-8. ISSN 0731-3985. LCCN QA75.5 .B766 1958. URL https://dl.

acm.org/doi/book/10.1145/
1458043.

Bruderer:2020:MAD

[Bru20]  Herbert Bruderer.  Mile-

stones in Analog and Digital Computing. Springer International Publishing, Cham, Switzerland, third edition, 2020. ISBN 3-030-40973-2, 3-030-40974-0 (e-book), 3-030-40975-

9. xxx + 2053 + 715 + 626 pp. LCCN QA76.17 .B78413

2020.  URL http://www.

springerlink.com/content/
978-3-030-40974-6.

Best:1936:EIV

[BRW36] J. E. Best, John Ashworth Ratcliffe, and Maurice Vincent Wilkes. Experimental investigations of very long wireless waves reflected from the ionosphere. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 156(889):614–633, September 1, 1936. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https:

//royalsocietypublishing.
org/doi/10.1098/rspa.1936.
0171.

Budden:1939:FIV

[BRW39] Kenneth George Budden, John Ashworth Ratcliffe, and Maurice Vincent Wilkes. Further investigations of very long waves reflected from the ionosphere. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 171(945):188–214, May 19, 1939. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https:

//royalsocietypublishing.
org/doi/10.1098/rspa.1939.
0061.

Boys:1954:EWF

[BS54] S. F. Boys and R. C. Sahni. Electronic wave functions XII. The evaluation of the general vectorcoupling coefficients by automatic computation. Philosophical Transactions of the Royal Society of London. Series A, Mathematical and Physical Sciences, 246(917):463–479, June 8, 1954. CODEN PTRMAD. ISSN 0080-4614 (print), 2054-0272 (electronic). URL https://

www.jstor.org/stable/91534.

Banerjee:1982:CR

[BS82] R. Banerjee and W. D. Shepherd. The Cambridge Ring. In Kenneth G. Beauchamp, editor, New Advances in Distributed Computer Systems: Proceedings of the NATO Advanced Study Institute held at Bonas, France, June 15–26, 1981, pages 223–

237. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 94-009-7760-3 (e-book), 94-009-7762-X (print). LCCN QA76.9.D5.

Beck:1996:CTN

[BS96] Robert E. Beck and Mary Lou Soffa, editors. CSC ’96: Proceedings of the 1996 ACM 24th Annual Conference on Computer Science, Philadelphia, Pennsylvania, USA, February 15–18, 1996. ACM Press, New York, NY, USA, 1996. ISBN 0-89791-828-2 (paperback). LCCN QA76 C7332 1996.

URL https://dl.acm.org/
doi/proceedings/10.1145/228329.

Bell:1982:CSP

[BSN82] C. Gordon Bell, Daniel P.

Siewiorek, and Allen Newell. Computer Structures: Principles and Examples. McGraw-Hill computer science series. Mc-Graw-Hill, New York, NY, USA, second. edition, 1982. ISBN 0-07-057302-6. xvi + 926 pp. LCCN QA76.9.A73 C65; TK7888.3 .B37 1982.

Bennett:1953:DCC

[BW53a] M. Bennett and M. V. Wilkes. Digital computation and the crystallographer. In Bowden [Bow53], pages 203–209. LCCN QA76.5 .B66. URL https://

                                        archive.org/details/FasterThanThought.

With a foreword by the Right Honourable the Earl of Halsbury.

Brooker:1953:FOE

[BW53b] R. A. Brooker and D. J. Wheeler. Floating operations on the EDSAC (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 7 (41):37–47, January 1953. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Buchholz:2000:WBC

[BW00a] W. Buchholz and M. Wilkes. Was Babbage caught in the act? IEEE Annals of the History of Computing, 22(4): 69–70, October/December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        85/19196/00887996.pdf.

Buchholz:2000:CQD

[BW00b] Werner Buchholz and Maurice Wilkes. Comments, queries, and debates: Was Babbage caught in the act? IEEE Annals of the History of Computing, 22(4):69–70, October/ December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2000/pdf/a4069.
pdf;   http://ieeexplore.
ieee.org/iel5/85/19196/00887996.
pdf.

Buchholz:1988:CQDc

[BWV^{+}88] Werner Buchholz, Maurice V.

Wilkes, Alfred W. Van Sinderen, C. J. Fern, Jr., and W. L. van der Poel. Comments, queries, and debate: Babbage and the Colossus; Babbage and Bowditch; Two Early European Computers; Early Dutch Computer. Annals of the History of Computing, 10(3):218–221, July/ September 1988. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/
an/books/an1988/pdf/a3218.
pdf;   http://www.computer.
org/annals/an1988/a3218abs.
htm.

Caminer:2003:BCL

[Cam03] David Tresman Caminer. Behind the curtain at LEO: a personal reminiscence. IEEE Annals of the History of Computing, 25(2):3–13, April/June 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        csdl.computer.org/dl/mags/
                                        an/2003/02/a2003.htm;   http:
                                        //csdl.computer.org/dl/mags/
                                        an/2003/02/a2003.pdf;   http:
                                        //csdl.computer.org/dl/mags/
                                        an/2003/02/a2003abs.htm.

Cochran:1955:UHS

[CD55] William Cochran and A. S. Douglas. The use of a high-speed digital computer for the direct determination of crystal structures. I. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 227(1171):486–500, February 8, 1955. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). See also part II [CD57].

Cochran:1957:UHS

[CD57] William Cochran and A. S. Douglas. The use of a high-speed digital computer for the direct determination of crystal structures. II. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 243(1233):281–288, December 24, 1957. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). See also part I [CD55].

Corbato:1963:CTS

[CDD^{+}63] F. J. Corbató, M. M. Daggett, R. C. Daley, R. J. Creasy, J. D. Hellwig, R. H. Orenstein, and L. K. Korn. The Compatible Time-Sharing System: a Programmer’s Guide. MIT Press, Cambridge, MA, USA, 1963. ISBN 0-262-03008-X. x + 96 pp. LCCN QA76 .M33. URL http://

www.bitsavers.org/pdf/mit/
ctss/CTSS_ProgrammersGuide.
pdf.

Ceruzzi:2003:HMC

[Cer03] Paul E. Ceruzzi. A History of Modern Computing. History of computing. MIT Press, Cambridge, MA, USA, second edition, 2003. ISBN 0-262-53203-4 (paperback). xi + 445 pp. LCCN QA76.17.C47 2003.

Ceruzzi:2013:SOP

[Cer13] Paul E. Ceruzzi. Scanning our past: Inventing the computer. Proceedings of the IEEE, 101(6):1503–1508, June 2013. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic). URL https://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6516706.

Campbell-Kelly:1980:FCP

[CK80a] Martin Campbell-Kelly. Foundations of Computer Programming in Britain 1945– 1955. Ph.D. thesis, Sunderland Polytechnic, Sunderland, UK, 1980. URL

https://www.proquest.com/
pqdtglobal/docview/1767363272.

Campbell-Kelly:1980:PEE

[CK80b] Martin Campbell-Kelly. Programming the EDSAC: Early programming activity at the University of Cambridge. Annals of the History of Computing, 2(1):7–36, January/ March 1980. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/
an/books/an1980/pdf/a1007.
pdf;   http://www.computer.
org/annals/an1980/a1007abs.

htm. Republished in [CK80b].

Campbell-Kelly:1992:ATE

[CK92a] Martin Campbell-Kelly. The Airy tape: An early chapter in the history of debugging. IEEE Annals of the History of Computing, 14(4):16–26, October/December 1992. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/a4016.
pdf;   http://www.computer.
org/annals/an1992/a4016abs.
htm.

Campbell-Kelly:1992:ICU

[CK92b] Martin Campbell-Kelly. Introduction: Computing at the University of Cambridge. IEEE Annals of the History of Computing, 14(4):8–??, October/ December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/a4008.

pdf. Special issue about computing at Cambridge under Maurice V. Wilkes.

Campbell-Kelly:1992:PEE

[CK92c] Martin Campbell-Kelly. Programming the EDSAC: early programming activity at the University of Cambridge. IEEE Annals of the History of Computing, 14(4):46–67, October/ December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/a4008.

pdf. Republication of [CK80b].

Campbell-Kelly:2011:PWW

[CK11]  Martin Campbell-Kelly.  In

praise of ’Wilkes, Wheeler, and Gill’. Communications of the ACM, 54(9):25–27, September 2011. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Campbell-Kelly:2014:SMV

[CK14] Martin Campbell-Kelly. Sir Maurice Vincent Wilkes. 26 June 1913–29 November 2010. Biographical Memoirs of Fellows of the Royal Society, 60:433–454, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:

//royalsocietypublishing.
org/doi/10.1098/rsbm.2013.
0020.

Campbell-Kelly:2019:EDC

[CK19a] Martin Campbell-Kelly. The evolution of digital computing practice on the Cambridge University EDSAC, 1949–1951. In Exploring the Early Digital [CK19b], pages 117–134. ISBN 3-030-02151-3 (e-book), 3-030-02152-1 (e-book), 3-030-02153-X (print). ISSN 2190-6831 (print), 2190-684X (electronic). LCCN QA76.17.

Campbell-Kelly:2019:EED

[CK19b] Martin Campbell-Kelly. Exploring the Early Digital. History of Computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2019. ISBN 3-030-02151-3 (e-book), 3-030-02152-1 (e-book), 3-030-02153-X (print). ISSN 2190-6831 (print), 2190-684X (electronic). xii + 203 + 30 pp. LCCN QA76.17.

Campbell-Kelly:1985:MSL

[CKW85] Martin Campbell-Kelly and Michael Roy Williams, editors. The Moore School Lectures: Theory and Techniques for Design of Electronic Digital Computers, volume 9 of Charles Babbage Institute reprint series for the history of computing. MIT Press, Cambridge, MA, USA, 1985. ISBN 0-262-03109-4. xxxiv + 568 + 2 pp. LCCN QA76.5 .M58 1985.

Copeland:2005:ATA

[Cop05] B. Jack Copeland, editor. Alan Turing’s Automatic Computing Engine: the master codebreaker’s struggle to build the modern computer. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-856593-3 (hardcover). xx + 553 pp. LCCN QA75 .A43 2005.

                                        URL http://ukcatalogue.
                                        oup.com/product/9780198565932.

do;   http://www.oxfordscholarship.
com/oso/public/content/maths/
9780198565932/toc.html.

Copeland:2006:MHC

[Cop06]  B. Jack Copeland.  The

modern history of computing. In Edward N. Zalta, editor, The Stanford Encyclopedia of Philosophy. Metaphysics Research Lab, Stanford University, Stanford, CA, USA, June

9, 2006. URL https://plato.
stanford.edu/entries/computing-
history/.

Copeland:2012:ATE

[Cop12] B. Jack Copeland, editor. Alan Turing’s Electronic Brain: the Struggle to Build the ACE, the World’s Fastest Computer. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-856593-3 (hardcover), 0-19-960915-2 (paperback), 0-19-164434-X (ebook). xxi + 553 pp. LCCN QA75 .C638 2012. UK £14.99.

Corbato:1958:BRP

[Cor58] Fernando J. Corbató. Book review: The Preparation of Programs for an Electronic Digital Computer, 2nd Revised Edition. By Maurice V. Wilkes, David J. Wheeler, Stanley Gill. 238 pp. Addison-Wesley Publishing Co., Reading, Mass., 1957. Physics Today, 11(8):28, August 1958. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Coghlan:2024:HPL

[CRB24] Brian Coghlan, Brian Randell, and Ralf Buelow. How Percy Ludgate’s 1909 paper (and IBM) helped thwart Konrad Zuse’s computer patent in 1960. IEEE Annals of the History of Computing, 46(3):20– 35, July/September 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Croarken:1990:ESC

[Cro90] Mary Croarken. Early Scientific Computing in Britain. Clarendon Press, Oxford, UK, 1990. ISBN 0-19-853748-4, 1-383-02607-6 (e-book). 160 + 8 pp. LCCN Q183.9 .C76 1990. URL http:

                                        //books.google.com/books?
                                        id=OW1QAAAAMAAJ;   http://
                                        catalog.hathitrust.org/api/
                                        volumes/oclc/20452712.html.

Croarken:1992:ECS

[Cro92] Mary G. Croarken. The emergence of computing science research and teaching at Cambridge, 1936–1949. IEEE Annals of the History of Computing, 14(4):10–15, October/ December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1992/pdf/a4010.
                                        pdf;   http://www.computer.
                                        org/annals/an1992/a4010abs.
                                        htm.

Cuff:2024:CMH

[Cuf24] Thomas Mark Cuff. Computer memories, a history, corrected, revised & expanded. Web document, August 25, 2024. URL

https://www.researchgate.
net/publication/355972939_
Computer_Memories_A_History_
Revision_3x.

Curley:2012:AIA

[Cur12a] Robert Curley. Architects of the Information Age, volume 1 of Computing and Connecting in the 21st Century. Britannica Educational Publishing, Chicago, IL, USA, 2012. ISBN 1-61530-717-6 (print), 1-61530-661-7. xviii + 156 pp. LCCN QA76.2.A2 A73 2012. 29.95 (NL).

Curley:2012:MW

[Cur12b] Robert Curley. Maurice Wilkes.

In Architects of the Information Age [Cur12a], chapter 15, page ?? ISBN 1-61530-717-6 (print), 1-61530-661-7. LCCN QA76.2.A2 A73 2012. 29.95 (NL).

Cohen:1999:MNH

[CWC99] I. Bernard Cohen, Gregory W.

Welch, and Robert V. D. Campbell, editors. Makin’ Numbers — Howard Aiken and the Computer. MIT Press, Cambridge, MA, USA, 1999. ISBN 0-262-03263-5 (hardcover), 0-262-27057-9, 0-585-13043-4 (ebook). xvii + 279 pp. LCCN QA76.2.A35 M35 1999.

URL https://dl.acm.org/
doi/book/10.5555/307688.

deBeauclair:2005:RMG

[dB05] Wilfried de Beauclair, editor. Rechnen mit Maschinen: Eine Bildgeschichte der Rechentechnik. (German) [Computing with Machines: a Visual History of Computing Technology]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 1-280-62377-2, 3-540-27433-2 (paperback), 3-540-27433-2 (e-book). 313 pp. LCCN QA76 .B325 2005; TK7885 .B438 2005.

Denning:1999:TBM

[Den99] Peter J. Denning, editor. Talking Back to the Machine. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 1-4612-2148-X. xxiv + 193 + 10 pp. LCCN QA75.5-76.95; QA76.9.C66. URL https://

link.springer.com/chapter/
10.1007/978-1-4612-2148-7_

11. Edited transcript of a lecture delivered at ACM97.

Douglas:1979:ASM

[Dou79] Sandy Douglas. Anecdotes: Some memories of EDSAC I: 1950–1952. Annals of the History of Computing, 1(2):98–99, 208, October/ December 1979. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/
an/books/an1979/pdf/a2098.
pdf;   http://www.computer.
org/annals/an1979/a2098abs.
htm.

Dwarkadas:2018:MWA

[Dwa18]  Sandhya Dwarkadas.  2018

Maurice Wilkes Award given to Gabriel Loh. IEEE Micro, 38(6):83–84, November/ December 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

URL https://www.computer.
org/csdl/mags/mi/2018/06/
08585397-abs.html.

ERA:1950:HSC

[Eng50] Engineering Research Associates, Inc. High-Speed Computing Devices. McGraw-Hill, New York, NY, USA, 1950. xiii + 451 pp. LCCN QA75 .E5. URL http://

www.bitsavers.org/pdf/era/
High_Speed_Computing_Devices_

1950.pdf. Contents of this volume were first assembled in the form of a report to the Office of Naval Research, prepared under a provision of contract N6-ONR-240, Task 1.

Everett:1952:WC

[Eve52] R. R. Everett. The Whirlwind I computer. In McPherson [McP52], pages 70–74. ISBN 1-4503-7851-X. LCCN QA75.5 .M374 1951. URL https://dl.

acm.org/doi/book/10.1145/
1434770;   https://ieeexplore.
ieee.org/xpl/conhome/5442625/
proceeding.

Fisher:1950:GFC

[Fis50] R. A. Fisher. Gene frequencies in a cline determined by selection and diffusion. Biometrics, 6(4):353–361, December 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.
jstor.org/stable/3001780.

Grant:1957:MSA

[GHR57] D. F. Grant, R. G. Howells, and D. Rogers. A method for the systematic application of sign relations. Acta Crystallographica, 10 (8):489–497, August 1957. CO-DEN ACCRA9. ISSN 0365-110x (print), 1600-8642 (electronic).

Gill:1951:DMP

[Gil51a] S. Gill. The diagnosis of mistakes in programmes on the ED-SAC. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 206(1087):538–554, May 1951. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https://

discovery.nationalarchives.
gov.uk/details/r/aeab6b06-
8815-4f40-8e24-0d2a4550428a.

Gill:1951:PSS

[Gil51b] S. Gill. A process for the step-by-step integration of differential equations in an automatic digital computing machine. Mathematical Proceedings of the Cambridge Philosophical Society, 47(1):96–108, January 1951. CODEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic). URL https://

discovery.nationalarchives.
gov.uk/details/r/aeab6b06-
8815-4f40-8e24-0d2a4550428a.

Gill:1953:AED

[Gil53] Stanley Gill. The Application of an Electronic Digital Computer to Problems in Mathematics and Physics. Ph.D. thesis, Faculty of Mathematics, University of Cambridge, Cambridge, UK, 1953. URL

https://www.proquest.com/
pqdtglobal/docview/1784058887/
16424FF6AECE4C50PQ/.

Gilchrist:1977:IPP

[Gil77] Bruce Gilchrist, editor. Information Processing, Proceedings of the 7th IFIP Congress 1977, Toronto, Canada, August 8–12, 1977. North-Holland, Amsterdam, The Netherlands, 1977. ISBN 0-7204-0755-9. LCCN

???? URL https://dblp.org/
db/conf/ifip/ifip1977.html.

Gill:1983:MW

[Gil83] S. Gill. Maurice Wilkes. In Anthony Ralston and Edwin D. Reilly, Jr., editors, Encyclopedia of Computer Science and Engineering, page ?? Van Nostrand Reinhold Co., New York, NY, USA, 1983. ISBN 0-442-24496-7. LCCN QA76.15.

Gilreath:2003:CAM

[GL03] William F. Gilreath and Phillip A.

Laplante. Computer Architecture: a Minimalist Perspective, volume 730 of The Springer International Series in Engineering and Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 1-4613-4980-X, 1-4615-0237-3 (e-book). ISSN 0893-3405. xv + 220 pp. LCCN QA76.9.A73 G53 2003; TK7895.M5.

Glass:1997:BRS

[Gla97] Robert L. Glass. In the Beginning: Recollections of Software Pioneers. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7999-9. vi + 318 pp. LCCN QA76.76.D47I517 1998. US$29.00. IEEE catalog number BP07999.

Gunther:1974:ICS

[GLL74] A. Günther, B. Levrat, and H. Lipps, editors. International Computing Symposium 1973: Proceedings of the International Computing Symposium 1973, Davos, Switzerland, 4–7 September 1973. North-Holland, Amsterdam, The Netherlands, 1974. ISBN 0-444-10596-4, 0-7204-2097-0. LCCN QA76 .I567 1973.

Griffith:1957:DPA

[GOB^{+}57] R. M. Griffith, J. A. Ovenstone, G. E. Barlow, J. H. Stapleton, and L. A. J. Malone, editors. Data Processing and Automatic Computing Machines: Proceedings of Conference Held at Weapons Research Establishment, Salisbury, SA, June 3rd–8th, 1957. Weapons Research Establishment, Salisbury, SA, Australia, 1957.

URL https://www.massey.ac.
nz/~rmclachl/DPACM/.

Gorman:1972:RGI

[Gor72] T. Gorman. Review of “The Growth of Interest in Microprogramming by M. V. Wilkes,” Computing Surveys, V1, #3, pp. 139–145 (September 1969). ACM SIG Micro Newsletter, 3 (3):58, October 1972. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/
1316540.1316545.

Grosch:1993:FDM

[Gro93] Herb Grosch. Forum: In defense of Maurice Wilkes (letter). Communications of the ACM, 36(5):16, May 1993. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hamblen:1958:RPP

[Ham58] John W. Hamblen. Recent publications: The Preparation of Programs for an Electronic Digital Computer, by M. V. Wilkes, D. J. Wheeler, and Stanley Gill. American Mathematical Monthly, 65(9):719–720, November 1958. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hartley:1999:EAC

[Har99] David Hartley. EDSAC 1 and after — a compilation of personal reminiscences. Cambridge University Web site, July 21, 1999. URL https:

//www.cl.cam.ac.uk/events/
EDSAC99/reminiscences/.

Hendry:1986:BRM

[Hen86] John Hendry. Book review: Memoirs of a Computer Pioneer by Maurice Wilkes. Isis, 77(3): 572–573, September 1986. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

                                        URL http://www.jstor.org/
                                        stable/231669.

Herbert:2024:RE

[Her24] Andrew Herbert. Reconstructing EDSAC. Web video (1h31m) and slides, April 25, 2024. URL

                                        https://leo-computers.org.
                                        uk/Videos/240426_EDSAC_(LEOSoc)
                                        .pdf;   https://leo-computers.
                                        org.uk/Videos/Andrew_Herbert_
                                        EDSAC_Rebuild_at_TNMoC.mp4.

Herbert:2004:CST

[HJ04] Andrew Herbert and Karen Spärck Jones, editors. Computer Systems: Theory, Technology, and Applications; a tribute to Roger Needham. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 0-387-20170-X. xvii + 338 pp. LCCN QA76.9.S88 H45 2004.

Hill:2000:RCA

[HJS00] Mark D. (Mark Donald) Hill, Norman P. (Norman Paul) Jouppi, and Gurindar Sohi, editors. Readings in Computer Architecture. Morgan Kaufmann Publishers, San Francisco, CA, USA, 2000. ISBN 1-55860-539-

8. xviii + 717 pp. LCCN QA76.9.A73 H55 2000. URL

http://books.elsevier.com/
bookscat/links/details.asp?
isbn=1558605398;   http://
www.loc.gov/catdir/description/
els033/99044480.html;   http:
//www.loc.gov/catdir/toc/
els033/99044480.html;   https:
//archive.org/details/readingsincomput0000hill/
page/n9/mode/2up?q=Slave+
memories;   https://shop.
elsevier.com/books/readings-
in-computer-architecture/
hill/978-0-08-057364-9.

Hartree:1948:DCM

[HNW^{+}48] D. R. Hartree, M. H. A. Newman, Maurice V. Wilkes, F. C. Williams, James H. Wilkinson, and A. D. Booth. A discussion on computing machines. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 195(1042):265– 287 (2 plates), December 22, 1948. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https://

ui.adsabs.harvard.edu/abs/
1948RSPSA.195..265H.

Hoffmann:2011:MWL

[Hof11]  Leah Hoffmann.  Maurice

Wilkes: the last pioneer. Communications of the ACM, 54 (2):19, February 2011. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

cacm.acm.org/news/maurice-
wilkes-the-last-pioneer/.

Hopper:1952:EC

[Hop52] Grace Murray Hopper. The education of a computer. In ????, editor, Proceedings of the ACM National Conference, May 1952, pages 243–249. ACM Press, New York, NY, USA, 1952.

Hartley:1981:CWA

[HT81] D. F. Hartley and The Editor. Correspondence: The Wilkes Award. The Computer Journal, 24(2):192, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_24/Issue_
                                        02/tiff/192.tif.

Haigh:2003:BOB

[HW03] Thomas Haigh and Maurice Wilkes. Biographies: Obituary: I. Bernard Cohen. IEEE Annals of the History of Computing, 25(4):89–92, October/December 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        csdl.computer.org/dl/mags/
                                        an/2003/04/a4089.htm;   http:
                                        //csdl.computer.org/dl/mags/
                                        an/2003/04/a4089.pdf.

Hoare:2004:RMN

[HW04] Tony Hoare and Maurice V. Wilkes. Roger Michael Needham CBE FREng. 9 February 1935–1 March 2003. Biographical Memoirs of Fellows of the Royal Society, 50:183–199, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:

                                        //royalsocietypublishing.

org/doi/epdf/10.1098/rsbm.
2004.0014.

IEEE:1992:TWW

[IEE92] IEEE, editor. Third Workshop on Workstation Operating Systems, April 23–24, 1992, Key Biscayne, Florida: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. IEEE catalog number 92TH0413-5.

Kendrew:1963:MSP

[Ken63] John C. Kendrew. Myoglobin and the structure of proteins: Crystallographic analysis and data-processing techniques reveal the molecular architecture. Science, 139(3561):1259– 1266, March 29, 1963. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Kidwell:1999:DRE

[Kid99] Peggy Aldrich Kidwell. Departments — reviews — electronic genie rescuing Prometheus; Forbes’ greatest technology stories; information ages; the mathematical theory of communication; LEO: The incredible story of the world’s first business; computer architects of the information society; Cryptologia; the supply of information technology workers in the US. IEEE Annals of the History of Computing, 21 (4):81–84, October/December 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        85/17389/00801538.pdf.

Kistermann:1992:CQDb

[Kis92]  F. W. Kistermann.  Com-

ments, queries, and debate: More on mechanical delay lines. IEEE Annals of the History of Computing, 14(4):7, October/December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

                                        URL http://dlib.computer.
                                        org/an/books/an1992/pdf/a4007.
                                        pdf.

Knuth:1976:EDP

[KT76] Donald E. Knuth and Luis Trabb Pardo. The early development of programming languages. Technical Report STAN-CS-76-562, Stanford University, Stanford, CA, USA, August 1976. i + 109 pp. URL

                                        https://bitsavers.org/pdf/
                                        stanford/cs_techReports/STAN-
                                        CS-76-562_EarlyDevelPgmgLang_
                                        Aug76.pdf.

Knuth:1977:EDP

[KT77] Donald E. Knuth and Luis Trabb Pardo. The early development of programming languages. In Belzer et al. [BHK77], pages 419–493. ISBN 0-8247-2251-5. LCCN QA76.15 .E5. Reprinted in [MHR80, pages 197–273].

Keckler:2014:ISC

[KT14] Stephen W. Keckler and Dean Tullsen. 2014 International Symposium on Computer Architecture Influential Paper Award; 2014 Maurice Wilkes Award given to Ravi Rajwar. IEEE Micro, 34(6):95–97, November/December 2014. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic).  URL http://www.
computer.org/csdl/mags/mi/
2014/06/mmi2014060095.html.

Land:1977:LFB

[Lan77] Frank Land. Leo, the first business computer: a personal experience. In In the Beginning: Recollections of Software Pioneers [Gla97], page ?? ISBN 0-8186-7999-9. LCCN QA76.76.D47I517 1998. US$29.00. IEEE catalog number BP07999.

Land:2000:FBC

[Lan00] F. Land. The first business computer: a case study in userdriven innovation. IEEE Annals of the History of Computing, 22(3):16–26, July/September 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
85/18655/00859523.pdf.

Laplante:1996:GPC

[Lap96] Phillip Laplante, editor. Great Papers in Computer Science. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-314-06365-X (paperback), 0-7803-1112-4 (hardcover). iv + 717 pp. LCCN QA76 .G686

1996. URL http://bit.csc.
lsu.edu/~chen/GreatPapers.
html.

Lavington:1980:CE

[Lav80a] Simon Hugh Lavington. The Cambridge EDSAC. In Early British Computers: the Story of Vintage Computers and the People Who Built Them [Lav80b], chapter 6, pages 31–

35. ISBN 0-932376-08-8 (paperback). LCCN TK7885.A5L38.

URL http://www.bitsavers.
org/pdf/dec/_Books/_Digital_
Press/Lavington_Early_British_
Computers_1980.pdf;   https:
//ed-thelen.org/comp-hist/
EarlyBritish.html.

Lavington:1980:EBC

[Lav80b] Simon Hugh Lavington. Early British Computers: the Story of Vintage Computers and the People Who Built Them. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1980. ISBN 0-932376-08-8 (paperback). 139 pp. LCCN TK7885.A5L38. URL http://

www.bitsavers.org/pdf/dec/
_Books/_Digital_Press/Lavington_
Early_British_Computers_1980.
pdf;   https://ed-thelen.
org/comp-hist/EarlyBritish.
html.

Lavington:2015:BHE

[Lav15] Simon Lavington. A brief history of early British computers. Web site, August 24, 2015. URL

https://ethw.org/A_Brief_
History_of_Early_British_
Computers;   https://ethw.
org/w/images/4/4f/Edsac1.
jpg.

Lee:1992:AId

[Lee92] J. A. N. Lee. About this issue. IEEE Annals of the History of Computing, 14(4):3–??, October/December 1992. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/a4003.
pdf.

Lee:1993:BWM

[Lee93]  J. A. N. Lee.  Birthday

wishes to Maurice V. Wilkes. IEEE Annals of the History of Computing, 15(3):5–??, July/ September 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1993/pdf/a3005.
pdf.

Lee:1995:MVW

[Lee95] J. A. N. Lee. Maurice Vincent Wilkes. Web site, 1995. URL

https://history.computer.
org/pioneers/pdfs/W/Wilkes.
pdf;   https://history.computer.
org/pioneers/wilkes.html.

Leslie:1983:ELA

[Les83] Ian Malcolm Leslie. Extending the Local Area Network.

Ph.D. thesis, Cambridge University, Cambridge, UK, May

17, 1983. URL https://www.
mathgenealogy.org/id.php?
id=137637;   https://www.
proquest.com/pqdtglobal/docview/
301416960/5371825BC8834A67PQ.

Lewis:2021:ICF

[Lew21] Harry R. Lewis, editor. Ideas That Created the Future: Classic Papers of Computer Science. MIT Press, Cambridge, MA, USA, 2021. ISBN 0-262-04530-

3. xxii + 495 pp. LCCN QA76 .I34 2021.

Leyden:2011:BPR

[Ley11] John Leyden. Bletchley Park to rebuild pioneering EDSAC computer: Will omit deadly 1940s mercury-based memory. The Register, ??(??):??, January 13, 2011. URL https:/

/www.theregister.com/2011/
01/13/edsac_rebuild_bletchley_
park/.

Liptay:1968:SAS

[Lip68] J. S. Liptay. Structural aspects of the SYSTEM/360 model 85, part II: The cache. IBM Systems Journal, 7(1):15–21, 1968. CO-DEN IBMSA7. ISSN 0018-8670.

Lennard-Jones:1939:DSD

[LJWB39] J. E. Lennard-Jones, M. V. Wilkes, and J. B. Bratt. The design of a small differential analyser. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 35(3):485–493, July 1939. CODEN PCPSA4.

ISSN 0008-1981. URL https:

//www.cambridge.org/core/
journals/mathematical-proceedings-
of-the-cambridge-philosophical-
society/article/abs/design-
of-a-small-differential-analyser/
E0A6F8EBECBE829B6AA80E8BFDF08436.

Lawson:1979:ESL

[LMW79] J. D. Lawson, J. Ll. Morris, and M. V. Wilkes. On the exact solution of linear constant coefficient parabolic partial differential equations using (1, 1) Padé approximations. Utilitas Mathematica, 16:189– 196, 1979. CODEN UTMADA. ISSN 0315-3681. URL https://

combinatorialpress.com/um/
vol16/.

Martonosi:2017:MWA

[Mar17] Margaret Martonosi. 2016 Maurice Wilkes Award given to Timothy Sherwood. IEEE Micro, 37 (2):104–105, March/April 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/mi/
2017/02/mmi2017020104.html.

Maxfield:1957:RPA

[Max57]  John E. Maxfield.  Recent

publications: Automatic Digital Computers, by M. V. Wilkes. American Mathematical Monthly, 64(3):205–206, March 1957. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

McLellan:2017:EEC

[McL17] Paul McLellan. ENIAC, ED-SAC and Colossus … and the Difference Engine. Web site, January 31, 2017. URL

                                        https://community.cadence.
                                        com/cadence_blogs_8/b/breakfast-
                                        bytes/posts/edsac.

McPherson:1952:RED

[McP52] J. C. McPherson, editor. Review of Electronic Digital Computers: Joint AIEE-IRE Computer Conference: Papers and Discussions Presented at the Joint AIEE-IRE Computer Conference, Philadelphia, Pa., December 10–12, 1951. ACM Press, New York, NY, USA, February 1952. ISBN 1-4503-7851-X. LCCN QA75.5 .M374 1951.

                                        URL https://dl.acm.org/
                                        doi/book/10.1145/1434770;
                                         https://ieeexplore.ieee.
                                        org/xpl/conhome/5442625/proceeding.

Metropolis:1980:HCT

[MHR80] Nicholas Metropolis, Jack Howlett, and Gian-Carlo Rota, editors. A History of Computing in the Twentieth Century: a Collection of Essays. Academic Press, New York, NY, USA, 1980. ISBN 0-12-491650-3, 1-4832-9668-7 (ebook). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Maulucci:1987:HEA

[MHW87] Ruth A. Maulucci, Harry D.

Huskey, and Maurice V. Wilkes. Happenings: ENIAC 40th Anniversary; Celebration at the Moore School; Eckert at the Computer Museum. Annals of the History of Computing, 9(2):199–202, April/ June 1987. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/
an/books/an1987/pdf/a2199.
pdf;   http://www.computer.
org/annals/an1987/a2199abs.
htm.

Miller:1951:LPN

[Mil51] J. C. P. Miller. Large prime numbers. Nature, 168(4280): 838, November 10, 1951. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Naur:1951:CSP

[Nau51a] P. Naur. Computation of special perturbations by an electronic calculator. Monthly Notices of the Royal Astronomical Society, 111(6):609–618, December 1951. CODEN MNRAA4. ISSN 1365-2966.

Naur:1951:MMP

[Nau51b] P. Naur. The motion of a minor planet or a comet. Edsac library specification sheet, University Mathematical Laboratory, Cambridge, UK, June 1951. 7 + 4 pp.

Naur:1951:VHS

[Nau51c] P. Naur. A vector-handling subroutine using call by name written for EDSAC in 1951. Report, University Mathematical Laboratory, Cambridge, UK, 1951. ???? pp.

Naur:1980:IED

[Nau80] Peter Naur. Impressions of the early days of programming. BIT, 20(4):414–425, December 1980. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0006-
3835&volume=20&issue=4&spage=
414.

Needham:1992:LDC

[Nee92] Roger M. Needham. Later developments at Cambridge: Titan, CAP, and the Cambridge Ring. IEEE Annals of the History of Computing, 14(4):57–58, October/December 1992. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/a4057.
pdf;   http://www.computer.
org/annals/an1992/a4057abs.
htm.

Needham:1982:CDC

[NH82] R. M. Needham and A. J. Herbert. The Cambridge Distributed Computing System. International computer science series. Addison-Wesley, Reading, MA, USA, 1982. ISBN 0-201-14092-6 (paperback). x + 170 pp. LCCN QA76.9.D5 N5 1982.

Needham:1974:DPM

[NW74] R. M. Needham and M. V.

Wilkes. Domains of protection and the management of processes. The Computer Journal, 17(2):117–120, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/17/2/117.full.pdf+
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
17/Issue_02/tiff/117.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/118.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/119.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_
02/tiff/120.tif.

Oettinger:1952:CPD

[Oet52] Anthony G. Oettinger. CXXIV.

Programming a digital computer to learn. Philosophical Magazine, 43(347):1243– 1263, December 1952. CODEN PHMAA4. ISSN 1941-5990.

ORegan:2018:ICC

[O’R18a] Gerard O’Regan. The Innovation in Computing Companion: a Compendium of Select, Pivotal Inventions. Springer International Publishing, Cham, Switzerland, 2018. ISBN 3-030-02618-3, 3-030-02619-1 (ebook), 3-030-02620-5. xxi + 300 pp. LCCN QA76.17.

ORegan:2018:LC

[O’R18b] Gerard O’Regan. LEO computers. In The Innovation in Computing Companion: a Compendium of Select, Pivotal Inventions [O’R18a], chapter 36, pages 175–178. ISBN 3-030-02618-3, 3-030-02619-1 (ebook), 3-030-02620-5. LCCN QA76.17.

ORegan:2018:MBM

[O’R18c] Gerard O’Regan. Manchester Baby and Mark 1 computers. In The Innovation in Computing Companion: a Compendium of Select, Pivotal Inventions [O’R18a], chapter 37, pages 179– 182. ISBN 3-030-02618-3, 3-030-02619-1 (e-book), 3-030-02620-5. LCCN QA76.17.

Orlowski:2010:BCP

[Orl10] Andrew Orlowski. British computer pioneer Wilkes reaches

                                        EOL: /etc/init.d/halt for

time sharing pioneer. The Register, ??(??):??, November 30, 2010. URL https:/

                                        /www.theregister.com/2010/
                                        11/30/maurice_wilkes_rip/.

Orgass:1965:WSCa

[OSWW65a] R. J. Orgass, H. Schorr, W. M. Waite, and M. V. Wilkes. WISP — a self-compiling list processing language. (for KDF9). Technical Report 36, Basser Computing Dept., University of Sydney, Sydney, NSW, Australia, 1965.

Orgass:1965:WSCb

[OSWW65b] R. J. Orgass, H. Schorr, W. M.

Waite, and M. V. Wilkes. WISP — a self-compiling list processing language. (for KDF9). Technical report, Yale Computer Centre, Yale University, New Haven, CT, USA, 1965.

Petard:1970:OSE

[Pet70] H. W. O. Petard. The optimum size for an establishment. Eureka, (33):19–??, ???? 1970. Communicated by M. V. Wilkes. Reprinted in [Pet92].

Petard:1992:OSE

[Pet92] H. W. O. Petard. The optimum size for an establishment. In R. L. Weber, editor, Science with a Smile, pages 96–?? Institute of Physics Publishing, Bristol, UK, 1992. ISBN 0-7503-0211-9. LCCN Q167 .S34 1992. Communicated by M. V. Wilkes. Reprint of [Pet70].

Pinkerton:1992:ICM

[PHL92] John M. M. Pinkerton, Derek Hemy, and Ernest H. Lenaerts. The influence of the Cambridge Mathematical Laboratory on the LEO project. IEEE Annals of the History of Computing, 14(4):41–48, October/ December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/a4041.

pdf;   http://www.computer.
org/annals/an1992/a4041abs.
htm.

Pollard:1953:COM

[PL53] B. W. Pollard and K. Lonsdale. The construction and operation of the Manchester University Computer. Journal of the Institution of Electrical Engineers, 1953(11):359, November 1953. CODEN JISEAL. ISSN 0368-2692 (print), 2054-0574 (electronic). URL https:

//digital-library.theiet.
org/docserver/fulltext/jiee-
2/1953/11/19530231.pdf.

Randell:1973:ODC

[Ran73] Brian Randell, editor. The Origins of Digital Computers: Selected Papers. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. ISBN 0-387-06169-X (New York), 3-540-06169-X. xvi + 464 pp. LCCN TK7888.3 .R36.

Randell:1975:ODC

[Ran75] Brian Randell, editor. The Origins of Digital Computers: Selected Papers. Texts and monographs in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1975. ISBN 0-387-07114-8, 3-540-07114-8, 3-642-96244-0, 3-642-96242-4 (ebook). xvi + 464 pp. LCCN ?TK7888.3 .R36 1975.

Randell:1982:ODC

[Ran82] Brian Randell, editor. The Origins of Digital Computers: Selected Papers. Texts and monographs in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1982. ISBN 0-387-11319-3, 3-540-11319-3. xvi + 580 pp. LCCN TK7885.A5 O741 1982.

Rojas:2000:FCH

[RH00] Raúl Rojas and Ulf Hashagen, editors. The First Computers: History and Architectures. History of computing. MIT Press, Cambridge, MA, USA, 2000. ISBN 0-262-18197-5 (hardcover), 0-585-35535-5 (electronic). xii + 457 pp. LCCN QA76.17 .F57 2000.

Richards:20xx:EIO

[Ricxx] Martin Richards. EDSAC initial orders and squares program. Web document, 20xx. URL

https://www.cl.cam.ac.uk/~
mr10/Edsac/edsacposter.pdf.

Ralston:1976:ECS

[RM76] Anthony Ralston and Chester L.

Meek, editors. Encyclopedia of Computer Science. Petrocelli/Charter, New York, NY, USA, 1976. ISBN 0-88405-321-0. xxviii + 1523 pp. LCCN QA76.15 .E48.

Ryle:1958:BLN

[Ryl58] Martin Ryle. Bakerian Lecture — the nature of the cosmic radio sources. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 248 (1254):289–308, November 1958. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

Scott:2014:MWA

[Sco14]  Steve Scott.  2013 Mau-

rice Wilkes Award given to Parthasarathy (Partha) Ranganathan. IEEE Micro, 34(1): 90–91, January/February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Shemer:1969:SAP

[SG69] J. E. Shemer and S. C. Gupta. A simplified analysis of processor “Look-Ahead” and simultaneous operation of a multi-module main memory. IEEE Transactions on Computers, C-18(1):64–71, January 1969. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671119.

Smith:2013:BPB

[Smi13] Tony Smith. Bletchley Park boffins start trailblazer ED-SAC computer rebuild: First replica parts go into production. The Register, ??(??):??, January 9, 2013. URL https:/

/www.theregister.com/2013/
01/09/bletchley_park_engineers_
begin_building_edsac_replica/

.

Stenstrom:2016:MWA

[Ste16] Per Stenstrom. 2015 Maurice Wilkes Award given to Christos Kozyrakis. IEEE Micro, 36(3):128–129, May/June 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/mi/
2016/03/mmi2016030128.html.

Strachey:1954:U

[Str54] Christopher S. Strachey. [unknown]. Ph.D. thesis, University of Cambridge, Cambridge, UK, 1954. URL https:/

/www.mathgenealogy.org/id.
php?id=75007.

Strachey:1959:PCT

[Str59a] Christopher Strachey. Programmecontrolled time sharing. Proceedings of the IEE — Part B: Electronic and Communication Engineering, 106(29):462, September 1959. CODEN PEE-BAE. ISSN 0369-8890 (print), 2054-0418 (electronic).

Strachey:1959:TSL

[Str59b] Christopher Strachey. Time sharing in large, fast computers. In ????, editor, Proceedings of the 1st International Conference on Information Processing, Paris, pages 336–341. UNESCO, Paris, France, 1959. URL

https://dblp.uni-trier.de/
db/conf/ifip/ifip1959.html.

Stanley:1950:TRG

[SW50] John Pearson Stanley and Maurice V. Wilkes. Table of the reciprocal of the gamma function for complex argument. Report, Computation Centre, University of Toronto, Toronto, ON, Canada, 1950. ???? pp.

Shinn:1952:GVG

[SW52] D. H. Shinn and H. A. Whale. Group velocities and group heights from the magneto-ionic theory. Journal of Atmospheric and Terrestrial Physics, 2(2): 85–105, January 1952. CODEN JASPF3. ISSN 0021-9169.

Strachey:1961:SPI

[SW61] Christopher Strachey and Maurce V. Wilkes. Some proposals for improving the efficiency of ALGOL 60. Communications of the ACM, 4(11): 488–491, November 1961. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stern:1980:CQD

[SW80] Nancy Stern and Maurice V. Wilkes. Comments, queries, and debate: Who “Invented” the first electronic digital computer?; Mauchly’s position on von Neumann’s role in drafting the EDVAC Report. Annals of the History of Computing, 2(4):375–377, October/ December 1980. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/
an/books/an1980/pdf/a4375.
pdf;   http://www.computer.
org/annals/an1980/a4375abs.
htm.

Swartzlander:1976:CDD

[Swa76] Earl E. Swartzlander, Jr. Computer Design Development: Principal Papers. Hayden Book Co., Rochelle Park, NJ, USA, 1976. ISBN 0-8104-5988-4. 310 pp. LCCN QA76.5 .C612565.

Taylor:1952:EEA

[Tay52] Norman H. Taylor. Evaluation of the engineering aspects of Whirlwind I. In McPherson [McP52], pages 75–

78. ISBN 1-4503-7851-X. LCCN QA75.5 .M374 1951.

URL https://dl.acm.org/
doi/book/10.1145/1434770;
https://ieeexplore.ieee.
org/xpl/conhome/5442625/proceeding.

Tucker:1967:MCS

[Tuc67] S. G. Tucker. Microprogram control for SYSTEM/360. IBM Systems Journal, 6(4):222–241, 1967. CODEN IBMSA7. ISSN 0018-8670.

Tropp:1987:ABE

[TW87] Henry S. Tropp and Maurice V.

Wilkes. Anecdotes: Babbage’s expectations for the Difference Engine. Annals of the History of Computing, 9(2):203–204, April/June 1987. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/
an/books/an1987/pdf/a2203.
pdf;   http://www.computer.
org/annals/an1987/a2203abs.
htm.

Uilks:1972:CSR

[Uil72] M. Uilks. Sistemy s razdeleniem vremeni. (Russian) [Timesharing systems]. Izdat. “Mir”, Moscow, USSR, 1972. 123 pp. Translated from the English by D. A. Korjagin, V. V. Smirnov and L. V. Uhov. Edited by È. Z. Ljubimskiĭ.

Uttley:1989:DPP

[UWE^{+}89] A. M. Uttley, M. V. Wilkes, W. S. Elliott, Brigadier G. H. Hinds, D. M. Mackay, and F. A. N. Hitch. Discussion of plans, projects, and general ideas. In Williams and Campbell-Kelly [WCK89], pages 147–158. ISBN 0-262-23136-0. LCCN QA76.17 .E171 1989.

vanderPoel:1982:ANM

[vdPWB82] W. L. van der Poel, Maurice V. Wilkes, and C. L. Baker. Anecdotes: Notes from the Mathematical Centre; First SHARE 709 Committee Meeting. Annals of the History of Computing, 4(3):275–283, July/ September 1982. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/
an/books/an1982/pdf/a3275.
pdf;   http://www.computer.
org/annals/an1982/a3275abs.

htm. Foreword by Morton I. Bernstein.

Verkley:1945:RCE

[VRW^{+}45] ?. Verkley, R. T. Ramm, M. V. Wilkes, R. T. P. Whipple, S. Devons, R. Bender, R. M. Whitmer, C. J. Carter, ?. Palmer, et al. Reports on certain establishments of the v Forschungsanstalt der Deutsche Reichspost (Research Institution of the German Post Office). Report, HMSO, London, UK, 1945. 44 pp.

Williams:1989:EBC

[WCK89] Michael Roy Williams and Martin Campbell-Kelly, editors. The Early British Computer Conferences, volume 14 of Charles Babbage Institute Reprint Series for the History of Computing. MIT Press, Cambridge, MA, USA, 1989. ISBN 0-262-23136-0. LCCN QA76.17 .E171 1989.

Wegner:1968:PLI

[Weg68] Peter Wegner. Programming Languages, Information Structures and Machine Organization. Ph.D. thesis, University of London, London, UK,

1968.  URL https://www.
mathgenealogy.org/id.php?
id=15166.

Wilkes:1982:MRJ

[WG82] Maurice V. Wilkes and I. J.

Good. Meetings in retrospect: J. G. Brainerd on the ENIAC; A Report on T. H. Flowers’s Lecture on Colossus. Annals of the History of Computing, 4(1):53–59, January/ March 1982. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/
an/books/an1982/pdf/a1053.
pdf;   http://www.computer.

                                        org/annals/an1982/a1053abs.
                                        htm.

Wilkes:1969:MSN

[WH69] M. V. Wilkes and D. F. Hartley. The management system — a new species of software? Datamation, 15(9):73–75, September 1969. CODEN DTMNAT. ISSN 0011-6963.

Wilkes:1997:CLS

[WH97] Maurice Wilkes and Andrew Hopper. The collapsed LAN: a solution to a bandwidth problem? ACM SIGARCH Computer Architecture News, 25(3): 1–5, June 1997. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Wheeler:1950:POI

[Whe50] D. J. Wheeler. Programme organization and initial orders for the EDSAC. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 202(1071):573–589, August 22, 1950. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

Wheeler:1951:ACE

[Whe51] David John Wheeler. Automatic Computing with EDSAC. Ph.D. thesis, Faculty of Mathematics, University of Cambridge, Cambridge, UK, 1951. URL

                                        https://www.proquest.com/
                                        pqdtglobal/docview/301253751/
                                         https://en.wikipedia.org/
                                        wiki/David_Wheeler_(computer_
                                        scientist).

Wheeler:1992:EPS

[Whe92a] David J. Wheeler. The ED-SAC programming systems. IEEE Annals of the History of Computing, 14(4):34–40, October/December 1992. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/a4034.
pdf;   http://www.computer.
org/annals/an1992/a4034abs.
htm.

Wheeler:1992:AE

[Whe92b] Joyce M. Wheeler. Applications of the EDSAC. IEEE Annals of the History of Computing, 14(4):27–33, October/ December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/a4027.
pdf;   http://www.computer.
org/annals/an1992/a4027abs.
htm.

Wilkes:1936:RVL

[Wil36] Maurice Vincent Wilkes. The reflection of very long wireless waves from the ionosphere. Ph.D. thesis, Cambridge University, Cambridge, UK, March 28,

1936. URL https://discover.
libraryhub.jisc.ac.uk/;
https://en.wikipedia.org/
wiki/Maurice_Wilkes;   https:
//www.britannica.com/biography/
Maurice-Wilkes;   https://
www.mathgenealogy.org/id.
php?id=143090;   https://
www.proquest.com/pqdtglobal/

                                        docview/301183158/DEC4F4836F2040F5PQ/
                                        4.

Wilkes:1939:TIC

[Wil39] M. V. Wilkes. Theoretical ionization curves for the E region. Proceedings of the Physical Society, London, 51(1):138– 146, January 1939. CODEN PPSOAU. ISSN 0959-5309 (print), 2051-2171 (electronic).

                                        URL https://iopscience.
                                        iop.org/article/10.1088/0959-
                                        5309/51/1/315.

Wilkes:1940:MSS

[Wil40a] M. V. Wilkes. A method of solving second order simultaneous linear differential equations using the Mallock machine. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 36(2):204–208, April 1940. CODEN PCPSA4. ISSN 0008-1981. URL https:

                                        //www.cambridge.org/core/
                                        journals/mathematical-proceedings-
                                        of-the-cambridge-philosophical-
                                        society/article/abs/method-
                                        of-solving-second-order-simultaneous-
                                        linear-differential-equations-
                                        using-the-mallock-machine/
                                        AE02EBB241E36AAB269BC519AFFBB2DE.

Wilkes:1940:RFI

[Wil40b] M. V. Wilkes. Region formation in the ionosphere according to an attachment theory. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 36(4):479–484, October 1940. CODEN PCPSA4.

ISSN 0008-1981. URL https:

//www.cambridge.org/core/
journals/mathematical-proceedings-
of-the-cambridge-philosophical-
society/article/abs/region-
formation-in-the-ionosphere-
according-to-an-attachment-
theory/7D7E75C572A887548273899415502C26.

Wilkes:1940:TRV

[Wil40c] Maurice Vincent Wilkes. The theory of reflexion of very long wireless waves from the ionosphere. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 175(961):143–163, April 10, 1940. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https:

//royalsocietypublishing.
org/doi/10.1098/rspa.1940.
0050.

Wilkes:1947:EHS

[Wil47a] M. V. Wilkes. The ENIAC — high-speed electronic calculating machine. Electronic Engineering, 19(230):104–108, 1947. CO-DEN ELEGAP. ISSN 0013-4902.

Wilkes:1947:ORV

[Wil47b] M. V. Wilkes. The oblique reflexion of very long wireless waves from the ionosphere. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 189(1016):130–147, March 27, 1947. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https:

//royalsocietypublishing.

                                        org/doi/10.1098/rspa.1947.
                                        0033.

Wilkes:1948:DPH

[Wil48] M. V. Wilkes. Design of a practical high-speed calculating machine. the EDSAC. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 195(1042):274–279, December 22, 1948. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https:

                                        //royalsocietypublishing.
                                        org/doi/10.1098/rspa.1948.
                                        0130.

Wilkes:1949:AO

[Wil49a] M. V. Wilkes. Atmospheric oscillations. Nature, 164(4163): 281, August 1949. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

                                        URL https://www.nature.
                                        com/articles/164281a0.

Wilkes:1949:ECM

[Wil49b]  M. V. Wilkes.  Electronic

calculating-machine development in Cambridge. Nature, 164(4170):557–558, October 1, 1949. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https:

                                        //www.nature.com/articles/
                                        164557a0.

Wilkes:1949:OEA

[Wil49c] M. V. Wilkes. Oscillations of the Earth’s Atmosphere. Cambridge monographs on physics. Cambridge University Press, Cambridge, UK, 1949. x + 76 pp. LCCN QC883 .W5.

Wilkes:1949:PDH

[Wil49d] M. V. Wilkes. Programme design for a high-speed automatic calculating machine. Journal of Scientific Instruments, 26(6):217–220, June 1949. CODEN JSINAY. ISSN 0368-4253, 0950-7671.

URL https://iopscience.
iop.org/article/10.1088/0950-
7671/26/6/318.

Wilkes:1949:PHS

[Wil49e] M. V. Wilkes. Progress in highspeed calculating machine design. Nature, 164(4165):341– 343, August 27, 1949. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL https://www.nature.
com/articles/164341a0.

Wilkes:1950:ACS

[Wil50a] M. V. Wilkes. Automatic computing. Nature, 166(4231):942– 943, December 2, 1950. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL https://www.nature.
com/articles/166942a0. Sum-

mer school report.

Wilkes:1950:RGL

[Wil50b] M. V. Wilkes. Rapport general (logical design). In ????, editor, Information processing: proceedings of the International Conference on Information Processing, page ?? UNESCO, Paris, France, 1950. ISSN 0074-3127. LCCN QA76 .I578; Q300 .F60.

Wilkes:1950:UEM

[Wil50c] M. V. Wilkes. The use of the EDSAC for mathematical computation. Applied Scientific Research B, 1:429–438, 1950. ISSN 0365-7140 (print), 1872-8065 (electronic).

Wilkes:1951:ACM

[Wil51a] M. V. Wilkes. Automatic calculating machines. Journal of the Royal Society of Arts, 100(4862):56–90, December 14, 1951. CODEN JRSAA4. ISSN 0035-9114. URL https://www.

                                        jstor.org/stable/41365298.

Three Cantor lectures, presented to the Royal Society of Arts, on 12, 19, and 26 November 1951.

Wilkes:1951:BWD

[Wil51b] M. V. Wilkes. The best way to design an automatic computing machine. In Williams and Kilburn [WK51], pages 16–18. LCCN ???? URL

                                        https://www.cs.princeton.
                                        edu/courses/archive/fall10/
                                        cos375/BestWay.pdf. Reprinted

in [Wil76a, ?, ?, ?, ?, ?].

Wilkes:1951:CMT

[Wil51c] M. V. Wilkes. Can machines think? Spectator, (6424):177– 178 (5–6 in issue), August 10, 1951. ISSN 0038-6952 (print), 2059-6499 (electronic). URL

                                        https://archive.spectator.
                                        co.uk/article/10th-august-
                                        1951/5/can-machines-think.

Wilkes:1951:TEA

[Wil51d] M. V. Wilkes. The thermal excitation of atmospheric oscillations. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 207(1090):358–370, July 6, 1951. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https:

//royalsocietypublishing.
org/doi/10.1098/rspa.1951.
0123.

Wilkes:1952:EC

[Wil52a] M. V. Wilkes. The EDSAC computer. In McPherson [McP52], pages 79–83. ISBN 1-4503-7851-X. LCCN QA75.5 .M374 1951.

URL https://dl.acm.org/
doi/book/10.1145/1434770;
https://ieeexplore.ieee.
org/xpl/conhome/5442625/proceeding.

Wilkes:1952:WWO

[Wil52b] M. V. Wilkes. World-wide oscillations in the Earth’s atmosphere. Quarterly Journal of the Royal Meteorological Society, 78(??):321–336, July 1952. CODEN QJRMAM. ISSN 0035-9009 (print), 1477-870X (electronic). URL https://

rmets.onlinelibrary.wiley.
com/doi/10.1002/qj.49707833703.

Symon’s Lecture.

Wilkes:1953:CMD

[Wil53a] M. V. Wilkes. Calculating machine development at Cambridge. In Bowden [Bow53], pages 130–134. LCCN QA76.5

.B66. URL https://archive.
org/details/FasterThanThought.

With a foreword by the Right Honourable the Earl of Halsbury.

Wilkes:1953:CMTa

[Wil53b] M. V. Wilkes. Can machines think? Discovery, 14(??): 151–??, ???? 1953. CODEN DISCAH. ISSN 0012-3625.

Wilkes:1953:CMTb

[Wil53c] M. V. Wilkes. Can machines think? Proceedings of the Institute of Radio Engineers, 41(10):1230–1234, 1953. CO-DEN PIREAE. ISSN 0096-8390 (print), 2162-6634 (electronic).

Wilkes:1953:E

[Wil53d] M. V. Wilkes. The EDSAC. In ????, editor, International Symposium on Automatic Digital Computation, National Physical Laboratory, March 1953, page ???? ????, ????, 1953. Reprinted in [WCK89, pages ??–??].

Wilkes:1953:EOE

[Wil53e] M. V. Wilkes. Expériences d’opérations effectuées avec l’E.D.S.A.C.. (French) [Operational experiences carried out with the E.D.S.A.C.]. In Joseph Pérès, editor, Les machines à calculer et la pensée humaine [Paris, 8–13 janvier 1951], number 37 in Colloques internationaux du Centre national de la recherche scientifique, pages 307–318. CNRS Paris, Paris, France, 1953.

Wilkes:1953:PAP

[Wil53f] M. V. Wilkes. Pure and applied programming. In Proceedings of the Association for Computing Machinery, Toronto, 1952, pages 121–124. Sauls Lithograph Co. (for the Association for Computing Machinery), Washington, DC, USA, 1953.

Wilkes:1953:UFA

[Wil53g] M. V. Wilkes. The use of a ‘floating address’ system for orders in an automatic digital computer. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 49(1):84–89, January 1953. CODEN PCPSA4. ISSN 0008-1981. URL https:

//www.cambridge.org/core/
journals/mathematical-proceedings-
of-the-cambridge-philosophical-
society/article/abs/use-of-
a-floating-address-system-
for-orders-in-an-automatic-
digital-computer/66DB2A4ACA578BB871B1B4A75352A6ED.

Wilkes:1954:TCG

[Wil54] M. V. Wilkes. A table of Chapman’s grazing incidence integral Ch(x, χ). Proceedings of the Physical Society, London, Section B, 67(4):304– 308, April 1954. CODEN PPSBAP. ISSN 0370-1301 (print), 1747-3837 (electronic).

URL https://iopscience.
iop.org/article/10.1088/0370-
1301/67/4/304.

Wilkes:1956:ADC

[Wil56a] M. V. Wilkes. Automatic Digital Computers. Methuen, London, UK, 1956. x + 305 pp. LCCN QA76 .W5 1956.

Wilkes:1956:DCH

[Wil56b] M. V. Wilkes. Digital computers and how they may help the engineer. Students’ Quarterly Journal, 27(105):26, September 1956. CODEN ???? ISSN 0039-2871 (print), 2053-7875 (electronic). URL https:

                                        //digital-library.theiet.
                                        org/content/journals/10.1049/
                                        sqj.1956.0048.

Wilkes:1956:NUA

[Wil56c] M. V. Wilkes. A note on the use of automativ [sic] adjustment of strip width in quadrature. In ????, editor, Report of a Conference on Electronic Digital Computers and Information Processing, Darmstadt, 25–27 October, 1955, pages 182–183. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1956. ISSN 0469-4325. LCCN QA76.5 .F3 1955. Offprint of Nachrichtentechnische Fachberichte (NTF), volume 4 (1956).

Wilkes:1956:SLA

[Wil56d] M. V. Wilkes. Solution of linear algebraic and differential equations by the long-division algorithm. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 52:758–763, 1956. CO-DEN PCPSA4. ISSN 0008-1981.

Wilkes:1956:TM

[Wil56e] M. V. Wilkes. Translation by machinery. Discovery, 17:83–??, ???? 1956. CODEN DISCAH. ISSN 0012-3625.

Wilkes:1957:ADC

[Wil57a] M. V. Wilkes. Automatic Digital Computers. Methuen & Co, London, UK, 1957. x + 305 pp. LCCN QA76.5 .W5 1957.

Wilkes:1957:CMT

[Wil57b] M. V. Wilkes. Can machines think? In ?. Pirelli, editor, ????, volume 10, pages 20–?? ????, ????, 1957.

Wilkes:1957:ISM

[Wil57c] M. V. Wilkes. Introduction to session on magnetic tape; input, output and auxiliary storage. Proceedings of the IEE — Part B: Radio and Electronic Engineering, 103(2S):331–332, April 1957. CODEN PEEBAE. ISSN 0369-8890 (print), 2054-0418 (electronic). URL https:

//digital-library.theiet.
org/content/journals/10.1049/
pi-b-1.1956.0068.

Wilkes:1957:SRN

[Wil57d]  M. V. Wilkes.  Some re-

marks on the numerical solution of ordinary differential equations. In Griffith et al. [GOB^{+}57], pages 108–1–108–6.

URL https://www.massey.ac.
nz/~rmclachl/DPACM/108%20-

%20Numerical%20solutions%
20of%20ODE.pdf.

Wilkes:1958:DCC

[Wil58a] M. V. Wilkes. Digital computers in control. Control, 1(??):204– ??, ???? 1958.

Wilkes:1958:OPD

[Wil58b] M. V. Wilkes. [Obituary:] Prof. D. R. Hartree, F.R.S. Nature, 181(4612):808, March 22, 1958. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Wilkes:1958:SDC

[Wil58c] M. V. Wilkes. The second decade of computer development. The Computer Journal, 1(3):98–105, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/1/3/98.full.pdf+html;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/010098.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/100.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/101.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/102.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/103.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_

03/tiff/104.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/105.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_
03/tiff/98.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_03/tiff/

99.tif. Presidential address.

Wilkes:1959:ADC

[Wil59a] M. V. Wilkes. Automatic Digital Computers. Dunod, Paris, France, 1959. ???? pp. LCCN ????

Wilkes:1959:ICI

[Wil59b] M. V. Wilkes. International conference on information processing. The Computer Bulletin, 3 (??):53–??, ???? 1959. CO-DEN COBUAH. ISSN 0010-4531 (print), 1464-357X (electronic).

Wilkes:1959:ILS

[Wil59c] M. V. Wilkes. Introductory lecture on special aspects of logical design. Proceedings of the IEE — Part B: Electronic and Communication Engineering, 106(29):462, September 1959. CODEN PEEBAE. ISSN 0369-8890 (print), 2054-0418 (electronic). URL https:

//digital-library.theiet.
org/content/journals/10.1049/
pi-b-2.1959.0310.

Wilkes:1959:M

[Wil59d] M. V. Wilkes. Microprogramming. In Broomall [Bro59], pages 18–20. ISBN 1-4503-7866-8. ISSN 0731-3985. LCCN QA75.5 .B766 1958. URL

                                        https://dl.acm.org/doi/10.
                                        1145/1458043.1458047.

Wilkes:1960:ACV

[Wil60] M. V. Wilkes. Avtomatičeskie cifrovye vyčislitel’nye mašiny. (Russian) [Automatic digital computers]. Staatlicher Unions-Verlag für Schiffbauindustrie, Leningrad, USSR, 1960. 356 pp.

Wilkes:1961:BRS

[Wil61a] M. V. Wilkes. Book review: Sampling Methods for Censuses and Surveys, Frank Yates, 1960 440 (London: Charles Griffin & Co. Ltd.). The Computer Journal, 3(4):219, 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.org/
                                        content/3/4/219.2.full.pdf+
                                        html.

Wilkes:1961:CTT

[Wil61b] M. V. Wilkes. Computer techniques of tomorrow. New Scientist, 11(??):805–??, ???? 1961. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Wilkes:1961:DTN

[Wil61c] M. V. Wilkes. Data transmission and the new outlook for the computer field. The Computer Journal, 4(1):1–9, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/040001.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/1.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
2.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/3.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/4.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/5.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/
6.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/7.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/8.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_
01/tiff/9.tif.

Wilkes:1961:DPS

[Wil61d] M. V. Wilkes. The design of practical self-repairing computer systems. Control, 4(??):86–??, ???? 1961.

Wilkes:1961:SRC

[Wil61e] M. V. Wilkes. Self-repairing computers. IRE Transactions on Electronic Computers, EC-10(1):93–94, March 1961. CODEN IRELAO. ISSN 0367-9950. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5219159.

Wilkes:1962:BRC

[Wil62a] M. V. Wilkes. Book review: Computer Handbook, by Harry D. Huskey, 1220 pages. (London: McGraw-Hill Publishing Co. Ltd.). The Computer Journal, 5(3):192, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.org/
                                        content/5/3/192.full.pdf+
                                        html.

Wilkes:1962:CTT

[Wil62b] M. V. Wilkes. Computer techniques of tomorrow. La Scuola in Agrone, 4(??):??, December 1962.

Wilkes:1962:DCD

[Wil62c] M. V. Wilkes. The digital computer in the drawing office. Control, 7(??):61–??, ???? 1962.

Wilkes:1962:PDC

[Wil62d]  M. V. Wilkes.  The place

of digital computers in control. In R. H. Macmillan, editor, Progress in Control Engineering, volume 1, page ?? Heywood, London, UK, 1962. LCCN TJ213 .P732.

Wilkes:1962:SLS

[Wil62e] M. V. Wilkes. The solar and luni-solar components of geomagnetic variation at San Fernando. Journal of Atmospheric and Terrestrial Physics, 24(2):

73–82, February 1962. CO-DEN JASPF3. ISSN 1364-6826 (print), 1879-1824 (elec-

tronic). URL https://www.
sciencedirect.com/science/
article/abs/pii/0021916962901873.

Wilkes:1963:ESC

[Wil63a] M. V. Wilkes. An experiment with a self-compiling compiler for a simple list-processing language. Technical memorandum 63/1, University of Cambridge Mathematical Laboratory, Cambridge, UK, 1963. 19 pp.

Wilkes:1963:OEA

[Wil63b] Maurice Vincent Wilkes. Oscillations of the Earth’s atmosphere with allowance for the variation of temperature with latitude. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 271 (1344):44–56, January 1, 1963. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https:

//royalsocietypublishing.
org/doi/10.1098/rspa.1963.
0004.

Wilkes:1964:CTS

[Wil64a] M. V. Wilkes. Constraint-type statements in programming languages. Communications of the ACM, 7(10):587–588, October 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilkes:1964:DMA

[Wil64b] M. V. Wilkes. The design of a multi-access computer system. In ????, editor, Proceedings of BCS/BIRE/IEE conference on ’The impact of users’ needs on the design of data processing systems, Edinburgh, 31 March–3 April, 1964, pages 87–?? ????, ????, 1964.

Wilkes:1964:LWT

[Wil64c] M. V. Wilkes. Lists and why they are useful. In E. A. Weiss, editor, ACM ’64: Proceedings of the 19th National Conference [of the ACM] Philadelphia, Pennsylvania, 25–27.08.1964, pages 61.1–61.5. ACM Press, New York, NY, USA, 1964. ISBN 1-4503-7918-4. LCCN QA75.5 .W457 1964. URL

                                        https://dl.acm.org/doi/10.
                                        1145/800257.808911.

Wilkes:1964:LTS

[Wil64d] M. V. Wilkes. On-line timeshared operation of a computer from a console. In Luigi Dadda and Umberto Pellegrini, editors, Automation and Instrumentation. Proceedings of the VIII International Convention, Milan, 19th–25th November, 1964, page ?? Pergamon Press, New York, NY, USA, 1964. LCCN TJ212 .I534 1967.

Wilkes:1964:WDC

[Wil64e] M. V. Wilkes. A world dominated by computers? New Scientist, ??(??):??, ???? 1964.

CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). Reprinted in [Wil65f].

Wilkes:1964:BDA

[Wil64f]  Maurice V. Wilkes.  Bab-

bage’s dream in action: [book reviews: Charles Babbage and His Calculating Engines, edited by Philip Morrison and Emily Morrison, New York: Dover, London: Constable, 1961]. Nature, 202(4933):631– 632, May 16, 1964. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/
nature/journal/v202/n4933/
pdf/202631a0.pdf.

Wilkes:1964:PUF

[Wil64g] Maurice V. Wilkes. A programmer’s utility filing system. The Computer Journal, 7(3):180–184, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/070180.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/180.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/181.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/182.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
03/tiff/183.tif;   http:/

                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_07/Issue_
                                        03/tiff/184.tif.

Wilkes:1964:ESC

[Wil64h] M. V. Wilkes, F.R.S. An experiment with a self-compiling compiler for a simple listprocessing language. In Richard Goodman, editor, Annual Review in Automatic Programming, volume 4 of International Tracts in Computer Science and Technology and Their Application, pages 1–48. Elsevier, Amsterdam, The Netherlands, 1964. URL https://www.

                                        sciencedirect.com/science/
                                        article/abs/pii/B9781483197807500055.

Wilkes:1965:ACC

[Wil65a] M. V. Wilkes. Abstracts of current computer literature. IEEE Transactions on Electronic Computers, EC-14(2):281–293, April 1965. CODEN IEECA8. ISSN 0367-7508. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=4038422.

Wilkes:1965:INA

[Wil65b] M. V. Wilkes. An Introduction to Numerical Analysis. Cambridge University Press, Cambridge, UK, 1965. 65 pp. LCCN ????

Wilkes:1965:LWT

[Wil65c] M. V. Wilkes. Lists and why they are useful. The Computer Journal, 7(4):278–281, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/7/4/278.full.pdf+
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
07/Issue_04/070278.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
07/Issue_04/tiff/278.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/279.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/280.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/281.tif.

Wilkes:1965:OTS

[Wil65d] M. V. Wilkes. Online time sharing — a very big step forward. Electronics and Power, 11(6):204–205, June 1965. CODEN ELPWAQ. ISSN 0013-5127 (print), 2053-7883 (electronic). URL https:

//digital-library.theiet.
org/content/journals/10.1049/
ep.1965.0166.

Wilkes:1965:SMD

[Wil65e] M. V. Wilkes. Slave memories and dynamic storage allocation. IEEE Transactions on Electronic Computers, EC-14(2):270–271, April 1965. CODEN IEECA8. ISSN 0367-7508. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4038419.

Wilkes:1965:WDC

[Wil65f] M. V. Wilkes. A world dominated by computers? In Nigel Calder, editor, The World in 1984, The complete New scientist series, page ?? Penguin Books, Harmondsworth, UK, 1965. LCCN Q171 .N5 1965. See [Wil64e].

Wilkes:1966:BRA

[Wil66a]  M. V. Wilkes.  Book re-

view: The Algebraic Eigenvalue Problem. The Computer Journal, 8(4):382, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.org/
                                        content/8/4/382.full.pdf+
                                        html.

Wilkes:1966:SIN

[Wil66b] M. V. Wilkes. A Short Introduction to Numerical Analysis. Cambridge University Press, Cambridge, UK, 1966. ISBN 0-511-66634-9, 0-521-09412-7 (paperback). iv + 76 pp. LCCN QA297 .W55.

Wilkes:1967:BRE

[Wil67a] M. V. Wilkes. [Book review]. Éléments de théorie des matrices carrées et rectangles en analyse numérique. (French) [Elements of theory of square and rectangle matrices in numerical analysis]. Electronics and Power, 13(5):176, May 1967. CODEN ELPWAQ. ISSN 0013-5127 (print), 2053-7883 (electronic). URL https:

//digital-library.theiet.
org/content/journals/10.1049/
ep.1967.0133.

Wilkes:1967:CMA

[Wil67b]  M. V. Wilkes.  The com-

ing of multiple-access computers. Journal of the Royal Aeronautical Society, 71(??):235–??, ???? 1967. CODEN JRAOAO. ISSN 0368-3931.

Wilkes:1967:CG

[Wil67c]  M. V. Wilkes.  Computer

graphics. Report 100, Centro Studi Calcolatrici Elettroniche (C.S.C.E.), Pisa, Italy, ???? 1967. ii + 18 + 4 pp. Fibonacci Lecture 1966.

Wilkes:1967:CO

[Wil67d] M. V. Wilkes. The computer outlook. Esso Magazine, 16(??): 22–??, ???? 1967. CODEN ES-MAAS. ISSN 0014-1011.

Wilkes:1967:DMA

[Wil67e]  M. V. Wilkes.  The de-

sign of multiple-access computer systems. The Computer Journal, 10(1):1–9, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/100001.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/1.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
2.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

Volume_10/Issue_01/tiff/3.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/4.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/5.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
6.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/7.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/8.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
01/tiff/9.tif.

Wilkes:1968:MAC

[Wil68a] M. V. Wilkes. Multiple access computer systems. Bulletin mathématique de la Société des Sciences Mathématiques de la République Socialiste de Roumanie, Nouvelle Série, 12(1):141–146, ???? 1968.

URL https://www.jstor.org/
stable/43679569.

Wilkes:1968:OIS

[Wil68b] M. V. Wilkes. The outer and inner syntax of a programming language. The Computer Journal, 11(3):260–263, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/110260.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/260.tif;   http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/261.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/262.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_
03/tiff/263.tif.

Wilkes:1968:TSCa

[Wil68c] M. V. (Maurice Vincent)

Wilkes. Time-sharing Computer Systems, volume 5 of Macdonald computer monographs. Macdonald and Co., London, UK, 1968. ISBN 0-356-02426-1, 0-444-19791-5. vii + 102 pp. LCCN QA76.5 .W523 1968b. Second edition 1972, Third edition 1975.

Wilkes:1968:TSCb

[Wil68d] M. V. (Maurice Vincent)

Wilkes. Time-sharing Computer Systems, volume 5 of Computer Monograph Series. American Elsevier Publishing Co., Inc., New York, NY, USA, 1968. vii + 102 pp. LCCN QA76.5 .W523.

Wilkes:1968:CTN

[Wil68e] Maurice V. Wilkes. Computers then and now. Journal of the ACM, 15(1):1–7, January 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wilkes:1969:BR

[Wil69a] M. V. Wilkes. Book review.

The Computer Journal, 12(4): 316, 1969. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/12/4/316.full.pdf+
html.

Wilkes:1969:GIM

[Wil69b] M. V. Wilkes. The growth of interest in microprogramming: a literature survey. ACM Computing Surveys, 1(3):139– 145, September 1969. CODEN CMSVAN. ISSN 0010-4892.

Wilkes:1969:MCS

[Wil69c] M. V. Wilkes. A model for core space allocation in a timesharing system. In Harrison W. Fuller, editor, AFIPS ’69 (Spring): Proceedings of the May 14–16, 1969, Spring Joint Computer Conference, pages 265–271. AFIPS Press, Montvale, NJ, USA, 1969. ISBN 1-4503-7902-8. LCCN ????

URL https://dl.acm.org/
doi/abs/10.1145/1476793.1476836.

Wilkes:1970:FFD

[Wil70] M. V. Wilkes. Forecasting the future: Development in computers. Long Range Planning, 2 (3):15–17, 1970. CODEN LR-PJA4, LRPDE4. ISSN 0024-6301 (print), 1873-1872 (elec-

tronic). URL https://www.
sciencedirect.com/science/
article/abs/pii/S0024630170800012.

Wilkes:1971:ATD

[Wil71a] M. V. Wilkes. Associative tabular data structures. ACM SIGPLAN Notices, 6(2):337– 345, February 1971. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See also revised version [Wil72a].

Wilkes:1971:ALAa

[Wil71b] M. V. Wilkes. Automatic load adjustment in time-sharing systems, 1971. Report on ACM SIGOPS Workshop on System Performance Evaluation, Harvard.

Wilkes:1971:ALAb

[Wil71c] M. V. Wilkes. Automatic load adjustment in time-sharing systems, 1971. Report on ACM SIGOPS Workshop on System Performance Evaluation, Seminaires IRIA, Le Chesnay, France.

Wilkes:1971:BCP

[Wil71d] M. V. Wilkes. Babbage as a computer pioneer. In ????, editor, The Babbage memorial meeting held 18th October 1971 at the Institution of Electrical Engineers building, Savoy Place, London WC2, report of proceedings, pages 1–18. British Computer Society and Royal Statistical Society, London, UK, 1971. LCCN ???? URL https:

//core.ac.uk/download/pdf/
82794809.pdf.

Wilkes:1971:CCS

[Wil71e] M. V. Wilkes. The changing computer scene 1947–71. In MC-25 Informatica Symposium, Tract 37. Math. Centre, Amsterdam, The Netherlands, 1971.

Wilkes:1971:HSI

[Wil71f] M. V. Wilkes. The hardware software interface. In IEEE, editor, Hardware software firmware trade-offs; conference digest: Fifth annual 1971 IEEE international Computer Society conference held at Boston, Sept. 22-24, 1971, pages 93–?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1971. LCCN QA76.5 .H3565. Abstract only.

Wilkes:1971:PSP

[Wil71g] M. V. Wilkes. The present status and potential of microprogramming. In IEEE Convention Record, page ?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1971. Abstract only.

Wilkes:1971:SMS

[Wil71h] M. V. Wilkes. Slave memories and segmentation. IEEE Transactions on Computers, C-20(6):674–675, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1671916.

Wilkes:1972:ATD

[Wil72a] M. V. Wilkes. Associative tabular data structures. International Journal of Computer and Information Sciences, 1(3):225– 233, September 1972. CODEN IJCIAH. ISSN 0091-7036. Revision of [Wil71a].

Wilkes:1972:BRA

[Wil72b] M. V. Wilkes. Book review:

Advances in computers, Vol. 10, Eds.: F. L. Alt and M. Rubinoff, Guest Ed.: W. Freiberger, Academic Press, New York, 1971. No. of pages: 313. Price: £6.75. Software—Practice and Experience, 2(2):191–192, April 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wilkes:1972:HPC

[Wil72c] M. V. Wilkes. Historical perspectives: Computer architecture. In Anonymous, editor, AFIPS ’72 (Fall, part II): Proceedings of the December 5– 7, 1972, [Anaheim, California], Fall Joint Computer Conference, part II, volume 42, pages 971–976. AFIPS Press, Montvale, NJ, USA, 1972. ISBN 1-4503-7913-3. LCCN ????

URL https://dl.acm.org/
doi/abs/10.1145/1480083.1480127.

Wilkes:1972:PID

[Wil72d]  M. V. Wilkes.  On pre-

serving the integrity of data bases. The Computer Journal, 15(3):191–194, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/15/3/191.full.pdf+
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_

15/Issue_03/150191.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_03/tiff/191.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/192.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/193.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_
03/tiff/194.tif;   https:/
/academic.oup.com/comjnl/
article/15/3/191/480522.

Wilkes:1972:SRV

[Wil72e] M. V. Wilkes. Sistemy s Razdeleniem Vremeni. (Russian) [Time-sharing Computer Systems]. Verlag “Mir”, Moscow, USSR, 1972. 124 pp. Translation from English to Russian by D. A. Korjagin, V. V. Smirnov and L. V. Uhov. Edited by E. Z. Ljubimskii.

Wilkes:1972:UWC

[Wil72f] M. V. Wilkes. The use of a writable control memory in a multiprogramming environment. In C. William Gear, editor, MICRO 5: Conference Record of the 5th Annual Workshop on Microprogramming, Urbana, Illinois, September 25–26, 1972, pages 62–65. ACM Press, New York, NY, USA, 1972. ISBN 1-4503-7369-0. LCCN QA76.6 .G437 1972. URL

https://dl.acm.org/doi/10.
1145/776378.776387.

Wilkes:1972:TSC

[Wil72g] M. V. (Maurice Vincent)

Wilkes. Time-sharing Computer Systems, volume 5 of Computer monographs. Macdonald and Co., London, UK, second edition, 1972. ISBN 0-444-19583-1 (Elsevier), 0-356-03985-4. ix + 149 pp. LCCN QA76.5 .W523 1972.

Wilkes:1973:CMA

[Wil73a] M. V. Wilkes. The Cambridge Multiple-Access System in retrospect. Software—Practice and Experience, 3(4):323–332, October/December 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wilkes:1973:DP

[Wil73b] M. V. Wilkes. The dynamics of paging. The Computer Journal, 16(1):4–9, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/160004.sgm.abs.html;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
01/tiff/4.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/
5.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/6.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/7.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_

                                        01/tiff/8.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_16/Issue_01/tiff/
                                        9.tif.

Wilkes:1974:ACP

[Wil74a] M. V. Wilkes. The art of computer programming, volume 3, sorting and searching. The Computer Journal, 17(4):324, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.org/
                                        content/17/4/324.full.pdf+
                                        html.

Wilkes:1974:PPF

[Wil74b] M. V. Wilkes. Past, present, and future in the computer field. In Günther et al. [GLL74], pages 615–620. ISBN 0-444-10596-4, 0-7204-2097-0. LCCN QA76 .I567 1973.

Wilkes:1974:STD

[Wil74c] M. V. Wilkes. Some thoughts on data bases. In D. Simpson and D. G. Rush, editors, Lectures in Advanced Programming 1973-4, pages 117–?? British Computer Society, ????, 1974. LCCN ????

Wilkes:1974:BRD

[Wil74d] Maurice V. Wilkes. Book review: Data base management, Infotech state of the art report 15, Ed. C. Boon and C. J. Bunyan, Infotech Information, Maidenhead, Berks, England, 1973. No. of pages: 665. Price: £40. Software—Practice and Experience, 4(3):301– 302, July/September 1974. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wilkes:1975:CUD

[Wil75a] M. V. Wilkes. Communication using a digital ring. In ????, editor, Proceedings of the Pacific Area Computer Communications Network (PACNET) Systems Symposium, Sendai, Japan, August 1975, pages 47– 55. ????, ????, August 1975. LCCN ????

Wilkes:1975:ECD

[Wil75b] M. V. Wilkes. Early computer developments at Cambridge: the EDSAC. The Radio and Electronic Engineer, 45(7):332–335, July 1975. CODEN RDEEA4. ISSN 0033-7722 (print), 2054-037X (electronic). URL https:

//digital-library.theiet.
org/content/journals/10.1049/
ree.1975.0063.

Wilkes:1975:EHI

[Wil75c] M. V. Wilkes. Early history of ionospheric investigations. Phil Trans. A, 280(??):54–??, ???? 1975.

Wilkes:1975:MPD

[Wil75d] M. V. Wilkes. Microprogramming principles and development. In ????, editor, Infotech State of the Art Report, volume 23, pages 163–?? ????, ????, 1975. LCCN ????

Wilkes:1975:TS

[Wil75e] M. V. Wilkes. Time sharing. In Philip B. Jordain and Michael Breslau, editors, Condensed Computer Encyclopedia, page ?? McGraw-Hill, New York, NY, USA, 1975. LCCN QA76.15 .J65.

Wilkes:1975:TSC

[Wil75f] M. V. (Maurice Vincent) Wilkes. Time-sharing Computer Systems. Computer monographs. Macdonald and Jane’s, London, UK, third edition, 1975. ISBN 0-444-19525-4 (American Elsevier), 0-356-08174-5. ii + 166 pp. LCCN QA76.53 .W54 1975.

Wilkes:1975:HBD

[Wil75g] Maurice V. Wilkes. How Babbage’s dream came true. Nature, 257(5527):541–544, October 16, 1975. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

                                        URL http://www.nature.com/
                                        nature/journal/v257/n5527/
                                        pdf/257541a0.pdf.

Wilkes:1976:BWD

[Wil76a] M. V. Wilkes. The best way to design an automatic calculating machine. In Computer Design Development: Principal Papers [Swa76], page ?? ISBN 0-8104-5988-4. LCCN QA76.5 .C612565.

Wilkes:1976:BRD

[Wil76b] M. V. Wilkes. Book review: Data base systems, Infotech State of the art Report, Infotech Information Ltd., Maidenhead, 1975. No. of pages: 722. Price: £60. Software—Practice and Experience, 6(4):582, October 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wilkes:1976:CB

[Wil76c] M. V. Wilkes. Charles Babbage.

In Ralston and Meek [RM76], pages 157–?? ISBN 0-88405-321-0. LCCN QA76.15 .E48.

Wilkes:1976:DA

[Wil76d] M. V. Wilkes. Differential analyser. In Ralston and Meek [RM76], pages 460–?? ISBN 0-88405-321-0. LCCN QA76.15 .E48.

Wilkes:1976:DCE

[Wil76e] M. V. Wilkes. Digital computer:

Early. In Ralston and Meek [RM76], pages 157–?? ISBN 0-88405-321-0. LCCN QA76.15 .E48.

Wilkes:1976:E

[Wil76f] M. V. Wilkes. EDSAC. In Ralston and Meek [RM76], pages 523–?? ISBN 0-88405-321-0. LCCN QA76.15 .E48.

Wilkes:1976:SES

[Wil76g] M. V. Wilkes. Software engineering and structured programming. IEEE Transactions on Software Engineering, SE-2(4):274–276, December 1976. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702383.

Wilkes:1976:UM

[Wil76h] M. V. Wilkes. Ultrasonic memory. In Ralston and Meek [RM76], pages 1439–?? ISBN 0-88405-321-0. LCCN QA76.15 .E48.

Wilkes:1977:AMMa

[Wil77a] M. V. Wilkes. The application of microcomputers to main frame design. In Anonymous, editor, Proceedings of the Conference on Computer Systems and Technology: Tuesday 29th to Thursday 31st March 1977, University of Sussex, volume 36 of I.E.R.E. conference proceedings, page ?? Institution of Electronic and Radio Engineers, London, UK, 1977. ISBN 0-903748-31-2 (paperback). LCCN QA76.5 .C619 1977.

Wilkes:1977:AMMb

[Wil77b] M. V. Wilkes. The application of microprocessors to main frame design. The Radio and Electronic Engineer, 47(7):302–304, July 1977. CODEN RDEEA4. ISSN 0033-7722 (print), 2054-037X (electronic). URL https:

//digital-library.theiet.
org/content/journals/10.1049/
ree.1977.0047.

Wilkes:1977:BCP

[Wil77c] M. V. Wilkes. Babbage as a computer pioneer. Historia Mathematica, 4(4):415– 440, November 1977. CO-DEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0315086077900799;
 https://core.ac.uk/download/
pdf/82794809.pdf.

Wilkes:1977:CF

[Wil77d] M. V. Wilkes. Computers in the future. In ????, editor, Proceedings, SEAS Anniversary Meeting 1977: September 19–23, 1977, Cambridge, England. Vol. I, page ?? ????, ????, 1977. ISSN 0254-6213. LCCN ???? Reprinted in [Wil78c].

Wilkes:1977:E

[Wil77e] M. V. Wilkes. The EDSAC. Report Com 90, National Physical Laboratory, Teddington, UK, 1977. Number 3 in a Series of Lectures on ‘The Pioneers of Computing’.

Wilkes:1977:FCS

[Wil77f] M. V. Wilkes. The future of computer systems. In ????, editor, Infotech State of the Art Report: Future Systems, pages 403–?? ????, ????, 1977. LCCN ????

Wilkes:1977:TYM

[Wil77g] M. V. Wilkes. Ten years and more of micro-programming. ACM SIG Micro Newsletter, 8 (4):11–13, December 1977. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/
1096518.1096519.

Wilkes:1977:BTC

[Wil77h] Maurice V. Wilkes. Beyond today’s computers. In Gilchrist [Gil77], pages 1–5. ISBN 0-7204-0755-9. LCCN ???? URL

                                        https://dblp.org/db/conf/
                                        ifip/ifip1977.html.

Wilkes:1978:BR

[Wil78a] M. V. Wilkes. Book review. The Computer Journal, 21(3): 245, 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.org/
                                        content/21/3/245.2.full.pdf+
                                        html.

Wilkes:1978:BRM

[Wil78b] M. V. Wilkes. [Book review]: Microprogrammable Computer Architectures. IEE Journal on Computers and Digital Techniques, 1(2):48, May 1978. CODEN IPETD3. ISSN 0140-1335 (print), 2053-8006 (electronic). URL https:

                                        //digital-library.theiet.
                                        org/content/journals/10.1049/
                                        ij-cdt.1978.0015.

Wilkes:1978:CF

[Wil78c] M. V. Wilkes. Computers in the future. Journal of the Computer Society of India, 8(??):3– ??, ???? 1978. CODEN JC-SIDS. ISSN 0379-5152. Reprint of [Wil77d].

Wilkes:1978:MMD

[Wil78d] M. V. Wilkes. Minicomputers, mainframes, and distributed systems. In Fred Maryanski, G. W. Gorsline, and Michael P. Deisenroth, editors, SIGMINI ’78: Proceedings of the first SIG-MINI symposium on Small systems, August 2–3, 1978, pages 1–6. ACM Press, New York, NY, USA, 1978. ISBN 1-4503-7870-6. LCCN ???? URL

https://dl.acm.org/doi/10.
1145/800218.806624.

Wilkes:1978:IPI

[Wil78e] Maurice V. Wilkes, editor.

ICSE ’78: Proceedings of the 3rd International Conference on Software Engineering, Atlanta, Georgia, USA, May 10– 12, 1978. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1978. LCCN QA76.6 .I553 1978.

Wilkes:1978:DCB

[Wil78f] M. V. Wilkes, F.R.S. Distributed computing in business data processing. ICL Technical Journal, (1):66– 70, November 1978. CO-DEN ITJOD7. ISSN 0142-

1557. URL https://vtda.org/
pubs/ICLTechnicalJournal/
ICL-Technical-Journal-v01i01.

pdf. Reprinted in [Wil79].

Wilkes:1979:DCB

[Wil79] M. V. Wilkes. Distributed computing in business data processing. Computer-Asia Software Guide, 6(??):32–??, ???? 1979. ISSN 8464-6320. Reprint of [Wil78f].

Wilkes:1980:Ca

[Wil80a]  M. V. Wilkes.  Computers

into the 1980s. Electronics and Power, 26(1):67–71, January 1980. CODEN ELPWAQ. ISSN 0013-5127 (print), 2053-7883 (electronic). URL https:

                                        //digital-library.theiet.
                                        org/content/journals/10.1049/
                                        ep.1980.0015.  See also

[Wil80b].

Wilkes:1980:Cb

[Wil80b] M. V. Wilkes. Computers into the 1980’s. In ????, editor, Infotech State of the Art Conference (March 1980), page ?? ????, ????, 1980. LCCN ???? See also [Wil80a].

Wilkes:1980:Cc

[Wil80c]  M. V. Wilkes.  Computers

into the 1980s. Electronics and Power, 26(1):67–71, January 1980. CODEN ELPWAQ. ISSN 0013-5127 (print), 2053-7883 (electronic).

Wilkes:1980:EPD

[Wil80d] M. V. Wilkes. Early programming developments in Cambridge. In Metropolis et al. [MHR80], pages 497–501. ISBN 0-12-491650-3, 1-4832-9668-7 (ebook). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Wilkes:1980:IWB

[Wil80e] M. V. Wilkes. The impact of wide-band local area communication systems on distributed computing. Computer, 13(9): 22–25, September 1980. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wilkes:1980:NHC

[Wil80f] M. V. Wilkes. A new hardware capability architecture. Operating Systems Review, 14(2):17– 20, April 1980. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wilkes:1981:BWD

[Wil81a] M. V. Wilkes. The best way to design an automatic computing machine. Microprocessing and Microprogramming, 8(3–5):141– 144, October/December 1981. CODEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (elec-

tronic). URL https://www.
sciencedirect.com/science/
article/abs/pii/0165607481900181.

Wilkes:1981:DCU

[Wil81b] M. V. Wilkes. The design of a control unit — reflections on reading Babbage’s notebooks. Annals of the History of Computing, 3(2):116–120, April/June 1981. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/
an/books/an1981/pdf/a2116.
pdf;   http://www.computer.
org/annals/an1981/a2116abs.
htm.

Wilkes:1982:HSMa

[Wil82a]  M. V. Wilkes.  Hardware

support for memory protection: Capability implementations. ACM SIGARCH Computer Architecture News, 10(2): 107–116, March 1982. CO-DEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Wilkes:1982:HSMb

[Wil82b] M. V. Wilkes. Hardware support for memory protection: Capability implementations. ACM SIGPLAN Notices, 17(4):107– 116, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wilkes:1982:KAP

[Wil82c] Maurice V. Wilkes. Keynote address — the processor instruction set. ACM SIG Micro Newsletter, 13(4):3–5, December 1982. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL

                                        https://dl.acm.org/doi/10.
                                        1145/1014194.800928.

Wilkes:1983:PMB

[Wil83a] M. V. Wilkes. Pray Mr. Babbage — a character study in dramatic form. Report, Computer Museum, Mountain View, CA, USA, 1983. See also [Wil91h].

Wilkes:1983:KJI

[Wil83b] Maurice V. Wilkes. Keeping jump instructions out of the pipeline of a RISC-like computer. ACM SIGARCH Computer Architecture News, 11(5):

5–7, December 1983. CO-DEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Wilkes:1983:SPS

[Wil83c] Maurice V. Wilkes. Size, power, and speed. ACM SIGARCH Computer Architecture News, 11(3):2–4, June 1983. CO-DEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE). Keynote address.

Wilkes:1984:HO

[Wil84a]  M. V. Wilkes.  Historical

overview. In ????, editor, AATS Invitation Symposium, page ?? Australian Academy of Technological Sciences and Engineering, Parkville, Victoria, Australia, 1984. ISSN 0157-7824. LCCN ????

Wilkes:1984:IRC

[Wil84b] M. V. Wilkes. Introduction to reprint of “Calculating Instruments and Machines” by D. R. Hartree. In ????, editor, Calculating machines: recent and prospective developments and their impact on mathematical physics; and, Calculating instruments and machines, volume 6 of Charles Babbage Institute Reprint Series, pages ix– ?? MIT Press, Cambridge, MA, USA, 1984. ISBN 0-262-08147-4. LCCN QA85 .H3 1984.

Wilkes:1984:PPF

[Wil84c] M. V. Wilkes. Past, present, and future in the computer field. IEE Proceedings, E:

Computers and Digital Techniques, 131(4):106–112, July 1984. CODEN IPETD3. ISSN 0143-7062 (print), 2053-7948 (electronic). URL https:

//digital-library.theiet.
org/content/journals/10.1049/
ip-e.1984.0020.

Wilkes:1984:RM

[Wil84d] Maurice Wilkes. Reflections on Maxwell. In Berger [Ber84], pages 191–192. ISBN 0-444-86707-4 (Elsevier). LCCN QC669 .J17 1984.

Wilkes:1984:SMP

[Wil84e] Maurice V. Wilkes. Security management and protection — a personal approach. The Computer Journal, 27(1):3–7, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/3.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/
4.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_27/Issue_01/tiff/5.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/6.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
01/tiff/7.tif.

Wilkes:1985:MCP

[Wil85] M. V. (Maurice Vincent) Wilkes. Memoirs of a Computer Pioneer. MIT Press series in the history of computing. MIT Press, Cambridge, MA, USA, 1985. ISBN 0-262-23122-0. viii + 240 pp. LCCN QA76.17 .W55 1985.

Wilkes:1986:GM

[Wil86a] M. V. Wilkes. The genesis of microprogramming. Annals of the History of Computing, 8(2):116–126, April/ June 1986. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/
an/books/an1986/pdf/a2116.
pdf;   http://www.computer.
org/annals/an1986/a2116abs.
htm.

Wilkes:1986:ESP

[Wil86b]  Maurice V. Wilkes.  EM-

PIRIC: a sketch of a programming language designed to facilitate a fine grain of protection. ACM SIGPLAN Notices, 21(8):16–27, August 1986. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Williams:1986:BRP

[Wil86c] Bernard O. Williams. Book review: The Preparation of Programs for an Electronic Digital Computer, with Special Reference to the “EDSAC” and the Use of a Library of Subroutines by Maurice V. Wilkes; David J. Wheeler; Stanley Gill. Isis, 77(1):157, March 1986. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/
stable/232546.

Wilkes:1987:RKE

[Wil87a] M. V. Wilkes. Recollections of King Edward’s School, Stourbridge (1921–31). The Black Countryman, 20(??):6–??, ???? 1987.

Wilkes:1987:CTN

[Wil87b] Maurice V. Wilkes. Computers then and now. In Ashenhurst [Ash87], page ?? ISBN 0-201-07794-9. LCCN QA76.24 .A33 1987. ACM Turing Award lecture.

Wilkes:1988:PFD

[Wil88] M. V. Wilkes. The past and future development of personal computers. Computer Standards and Interfaces, 8(1):5–7, 1988. CODEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (elec-

                                        tronic). URL https://www.
                                        sciencedirect.com/science/
                                        article/abs/pii/0920548988900694.

Wilkes:1989:BWD

[Wil89a] M. V. Wilkes. The best way to design an automatic calculating machine. In Williams and Campbell-Kelly [WCK89], pages 182–184. ISBN 0-262-23136-0. LCCN QA76.17 .E171 1989.

Wilkes:1989:Ea

[Wil89b] M. V. Wilkes. The E.D.S.A.C. In Williams and Campbell-Kelly [WCK89], pages 16–20. ISBN 0-262-23136-0. LCCN QA76.17 .E171 1989.

Wilkes:1989:Eb

[Wil89c] M. V. Wilkes. The E.D.S.A.C.

In Williams and Campbell-Kelly [WCK89], pages 229–231. ISBN 0-262-23136-0. LCCN QA76.17 .E171 1989.

Wilkes:1990:NEF

[Wil90a]  Maurice Wilkes.  Networks,

email and fax. Communications of the ACM, 33(6):631– 633, June 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilkes:1990:BF

[Wil90b] Maurice V. Wilkes. The bandwidth famine. Communications of the ACM, 33(8):19– 21, August 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-0782/
79174.html.

Wilkes:1990:BAM

[Wil90c] Maurice V. Wilkes. The British Association Mathematical Tables Committee. Historia Mathematica, 17(2):152, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/031508609090051E.

Wilkes:1990:CSB

[Wil90d] Maurice V. Wilkes. Computer security in the business world. Communications of the ACM, 33(4):399–401, 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilkes:1990:CBS

[Wil90e] Maurice V. Wilkes. Computers before silicon: design decisions on EDSAC. IEE Review, 36(11):429–431, December 1990. CODEN IEREEF. ISSN 0953-5683 (print), 1741-0495 (electronic). URL https:

//digital-library.theiet.
org/content/journals/10.1049/
ir_19900181.

Wilkes:1990:CPA

[Wil90f]  Maurice V. Wilkes.  Com-

puting perspectives: It’s all software, now. Communications of the ACM, 33(10):19– 21, October 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-0782/
84539.html.

Wilkes:1990:CPR

[Wil90g] Maurice V. Wilkes. Computing perspectives: the rise of the VLSI processor. Communications of the ACM, 33(12):16–17, 44, December 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-0782/
96286.html.

Wilkes:1990:PCP

[Wil90h] Maurice V. Wilkes. A perspective on computer progress in the last five years. Computing and Control Engineering Journal, 5(1):201–208, September 1990. CODEN CCEJEL. ISSN 0956-3385 (print), 1741-0460 (electronic). URL https:

//digital-library.theiet.
org/content/journals/10.1049/
cce_19900060.

Wilkes:1990:HPB

[Wil90i] M. V. Wilkes, F.R.S. Herschel, Peacock, Babbage and the development of the Cambridge curriculum. Notes and Records of the Royal Society of London, 44(2):205–219, July 1990. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic).  URL http://www.
jstor.org/stable/531607.

Wilkes:1991:EMV

[Wil91a] M. V. Wilkes. Editorial — the maturing of VLSI — and its consequences for the computer industry. The Computer Journal, 34(3):193–194, June 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/34/3/193.full.pdf+
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
34/Issue_03/tiff/193.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_
03/tiff/194.tif.

Wilkes:1991:BCP

[Wil91b]  Maurice V. Wilkes.  Bab-

bage as a computer pioneer. British Computer Society and the Royal Statistical Society, ?? (??):??, 1991.

Wilkes:1991:BEH

[Wil91c]  Maurice V. Wilkes.  Bab-

bage’s expectations for his Engines. Annals of the History of Computing, 13(2):141–145, April/June 1991. CODEN AH-COE5. ISSN 0164-1239. URL

                                        http://dlib.computer.org/
                                        an/books/an1991/pdf/a2141.
                                        pdf;   http://www.computer.
                                        org/annals/an1991/a2141abs.
                                        htm.

Wilkes:1991:CBH

[Wil91d] Maurice V. Wilkes. Charles Babbage and his calculating engines. IEE Review, 38 (2):71, February 1991. CO-DEN IEREEF. ISSN 0953-5683 (print), 1741-0495 (electronic). URL https://

                                        digital-library.theiet.org/
                                        content/journals/10.1049/
                                        ir_19920030.

Wilkes:1991:CPC

[Wil91e] Maurice V. Wilkes. Computing perspectives: Could it have been predicted? Communications of the ACM, 34(10):19– 21, October 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org/
                                        pubs/toc/Abstracts/0001-0782/
                                        125288.html.

Wilkes:1991:CPR

[Wil91f]  Maurice V. Wilkes.  Com-

puting perspectives: Revisiting computer security in the business world. Communications of the ACM, 34(8):19– 22, August 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-0782/
108519.html.

Wilkes:1991:CPS

[Wil91g]  Maurice V. Wilkes.  Com-

puting perspectives: software and the programmer. Communications of the ACM, 34 (5):23–24, May 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-0782/
103169.html.

Wilkes:1991:PMB

[Wil91h]  Maurice V. Wilkes.  Pray

Mr. Babbage — a character study in dramatic form. Annals of the History of Computing, 13(2):147–154, April/ June 1991. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/
an/books/an1991/pdf/a2147.
pdf;   http://www.computer.
org/annals/an1991/a2147abs.

htm. See also [Wil83a].

Wilkes:1992:CNA

[Wil92a] M. V. Wilkes. The case for a new approach to operating systems for personal computers and work stations. In IEEE [IEE92], pages 164–167. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

/ieeexplore.ieee.org/iel2/

                                        891/6831/00275668.pdf. IEEE

catalog number 92TH0413-5.

Wilkes:1992:PRC

[Wil92b] M. V. Wilkes. Progress and research in the computer industry. Philosophical Transactions of the Royal Society of London: Physical Sciences and Engineering, 334(1633): 173–184, January 15, 1992. ISSN 0962-8428 (print), 2054-0299 (electronic). URL https:

                                        //royalsocietypublishing.
                                        org/doi/10.1098/rsta.1991.

0006. Clifford Paterson Lecture 1990.

Wilkes:1992:WJK

[Wil92c] M. V. Wilkes. Wirukusu jiden : Konpyūta no paionia no kaisō. Maruzen, Tokyo, Japan, 1992. ISBN 4-621-03743-9. x + 302 pp. LCCN ????

                                        URL https://ci.nii.ac.jp/
                                        ncid/BN08137795. Translation

to Japanese by Nobue Nakamura and Akira Nakamura of [Wil85].

Wilkes:1992:CBG

[Wil92d] Maurice V. Wilkes. Charles Babbage — the great uncle of computing? Communications of the ACM, 35(3):15–21, 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilkes:1992:CPA

[Wil92e]  Maurice V. Wilkes.  Com-

puting perspectives: Artificial intelligence as the year 2000 approaches. Communications of the ACM, 35(8):17– 23, August 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-0782/
135237.html.

Wilkes:1992:CPC

[Wil92f]  Maurice V. Wilkes.  Com-

puting perspectives: Charles Babbage: The Great Uncle of computing? Communications of the ACM, 35(3):15– 16, 21, March 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-0782/
214839.html.

Wilkes:1992:CPL

[Wil92g]  Maurice V. Wilkes.  Com-

puting perspectives. the longterm future of operating systems. Communications of the ACM, 35(11):23–24, 112, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-0782/
138855.html.

Wilkes:1992:E

[Wil92h] Maurice V. Wilkes. EDSAC 2. IEEE Annals of the History of Computing, 14(4):49–56, October/December 1992. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an1992/pdf/a4049.

                                        pdf;   http://www.computer.
                                        org/annals/an1992/a4049abs.
                                        htm.

Wilkes:1993:CPF

[Wil93a] Maurice V. Wilkes. Computing perspectives: From Fortran and Algol to objectoriented languages. Communications of the ACM, 36(7): 21–23, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org/
                                        pubs/toc/Abstracts/0001-0782/
                                        159553.html.

Wilkes:1993:CPL

[Wil93b] Maurice V. Wilkes. Computing perspectives: Light amplifiers and solitons. Communications of the ACM, 36(2):21– 23, February 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org/
                                        pubs/toc/Abstracts/0001-0782/
                                        151221.html.

Wilkes:1994:OSC

[Wil94a] Maurice Wilkes. Operating systems in a changing world. Operating Systems Review, 28(2): 9–21, April 1994. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wilkes:1994:EOD

[Wil94b] Maurice V. Wilkes. The ED-SAC: Origins and design decisions. In Klaus Brunnstein and Eckart Raubold, editors, Applications and Impacts, Information Processing ’94, Volume 2, Proceedings of the IFIP 13th World Computer Congress, Hamburg, Germany, 28 August– 2 September, 1994, volume A-52 of IFIP Transactions, pages 301–306. North-Holland, Amsterdam, The Netherlands, 1994. ISBN 0-444-81987-8 (paperback). LCCN QA76.9.A43 I34 1992.

Wilkes:1995:CP

[Wil95a] M. V. Wilkes. Computing Perspectives. Morgan Kaufmann Publishers, San Francisco, CA, USA, 1995. ISBN 0-08-050268-7, 1-322-46561-4, 1-55860-317-4. x + 207 pp. LCCN QA76 .W5 1995; QA76.17 .C667 1995.

Wilkes:1995:BRU

[Wil95b]  Maurice V. Wilkes.  Book

review: The ubiquitous Mr Babbage: Passages from the Life of a Philosopher [Charles Babbage, Passages from the Life of a Philosopher (ed. M. Campbell-Kelly). London. Pickering and Chatto, 1994. Pp. 392, £12.95. ISBN 1-85196-040-6]. Notes and Records of the Royal Society of London, 49 (2):341–342, July 1995. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic).  URL http://www.
jstor.org/stable/532024.

Wilkes:1995:MWC

[Wil95c] Maurice V. Wilkes. The memory wall and the CMOS endpoint. ACM SIGARCH Computer Architecture News, 23(4):

4–6, September 1995. CO-DEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Wilkes:1995:TPE

[Wil95d] Maurice V. Wilkes. A tribute to Presper Eckert. Communications of the ACM, 38(9):20– 22, September 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org/
                                        pubs/toc/Abstracts/0001-0782/
                                        223253.html.

Wilkes:1996:BWD

[Wil96a]  M. V. Wilkes.  The best

way to design an automatic calculating machine. In Laplante [Lap96], pages 279– 286. ISBN 0-314-06365-X (paperback), 0-7803-1112-4 (hardcover). LCCN QA76 .G686

                                        1996. URL http://bit.csc.
                                        lsu.edu/~chen/GreatPapers.
                                        html.

Wilkes:1996:CEP

[Wil96b] M. V. Wilkes. The CMOS end-point and related topics in computing. Computing and Control Engineering Journal, 7(2):101–106, April 1996. CODEN CCEJEL. ISSN 0956-3385 (print), 1741-0460 (electronic). URL https:

                                        //digital-library.theiet.
                                        org/content/journals/10.1049/
                                        cce_19960207.

Wilkes:1996:CTN

[Wil96c]  M. V. Wilkes.  Computers

then and now — part 2. In Beck and Soffa [BS96], pages 115–119. ISBN 0-89791-828-2 (paperback). LCCN QA76 C7332 1996. URL https:

//dl.acm.org/doi/10.1145/
228329.228342.

Wilkes:1996:TDB

[Wil96d] M. V. Wilkes. Technology drive:

[book review]: Computer: A History of the Information Machine. By Martin Campbell-Kelly and William Aspray. BasicBooks: 1996. Pp. 342. Nature, 383(6599):405, October 3, 1996. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https:

//www.nature.com/articles/
383405a0.

Wilkes:1996:UDI

[Wil96e] M. V. Wilkes. User-driven innovation: the world’s first business computer. IEE Review, 42(4): 165–, 1996. CODEN IEREEF. ISSN 0953-5683 (print), 1741-0495 (electronic).

Wilkes:1997:AE

[Wil97a]  M. V. Wilkes.  Arithmetic

on the EDSAC. IEEE Annals of the History of Computing, 19(1):13–15, January/ March 1997. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:/

/ieeexplore.ieee.org/iel4/
85/12228/00560726.pdf.

Wilkes:1997:CPD

[Wil97b]  M. V. Wilkes.  Challenges

for processor designers. Computer, 30(1):10–??, January 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

//ieeexplore.ieee.org/xpl/
tocresult.jsp?isnumber=12229&
punumber=2.

Wilkes:1997:CCR

[Wil97c] M. V. Wilkes. Contribution to ‘Cockcroft remembered’, a symposium held in Churchill College on 29 May 1997. Churchill Magazine, ??(??):24–??, ???? 1997.

Wilkes:1997:SEA

[Wil97d] M. V. Wilkes. Sir Edward Appleton and early ionosphere research. Notes and Records of the Royal Society of London, 51(2): 281–290, July 22, 1997. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic).  URL http://www.
jstor.org/stable/531992.

Wilkes:1998:CWS

[Wil98a] M. V. Wilkes. CMOS workstations and servers — how far can evolution and innovation take us. In IEEE, editor, PACT ’98: Proceedings of the 1998 International Conference on Parallel Architectures and Compilation Techniques, Paris, France, October 12–18, 1998, pages 1–9. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-8591-3. LCCN QA76.58 .I5445 1999.

URL https://www.cl.cam.ac.
uk/archive/mvw1/pact98.pdf.

Wilkes:1998:SP

[Wil98b] M. V. Wilkes. A swing of the pendulum? Microprocessor Report, 12(8):1–2, June 22, 1998. ISSN 0899-9341 (print), 1943-5231 (electronic). URL https:

//www.ardent-tool.com/CPU/
docs/MPR/19980622/1208vp.
pdf.

Wilkes:1998:RAE

[Wil98c] Maurice V. Wilkes. A revisionist account of early language development. Computer, 31(4):22– 25, April 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1998/pdf/r4022.
pdf;   http://www.computer.
org/computer/co1998/r4022abs.
htm.

Wilkes:1999:HCS

[Wil99a] M. V. Wilkes. A half century of surprises. In Denning [Den99], pages 111–116. ISBN 1-4612-2148-X. LCCN QA75.5-76.95; QA76.9.C66. URL https://

link.springer.com/chapter/
10.1007/978-1-4612-2148-7_

11. Edited transcript of a lecture delivered at ACM97.

Wilkes:1999:HSS

[Wil99b] M. V. Wilkes. Historical studies in science and technology and the uses to which they can be put. Notes and Records of the Royal Society of London, 53(1): 3–10, January 22, 1999. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/531925.

Wilkes:1999:U

[Wil99c] M. V. Wilkes. A view from overseas. In Cohen et al. [CWC99], pages 215–218. ISBN 0-262-03263-5 (hardcover), 0-262-27057-9, 0-585-13043-4 (ebook). LCCN QA76.2.A35 M35 1999. URL https:

                                        //dl.acm.org/doi/10.5555/
                                        307688.307703.

Wilkes:2000:CBC

[Wil00a] M. V. Wilkes. The cogwheel brain: Charles Babbage and the quest to build the first computer [book review]. IEEE Annals of the History of Computing, 22(4):62–63, October/December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        85/19196/00887990.pdf.

Wilkes:2000:REV

[Wil00b] M. V. Wilkes. Recollections of early vacuum tube circuits. Computer Resurrection (BCS Computer Conservation Society), (24):7–9, Autumn 2000. ISSN 0958-7403.

                                        URL http://www.cs.man.ac.
                                        uk/CCS/res/res24.htm#c;
                                        https://waldorf.cs.man.ac.
                                        uk/ftp/pub/CCS-Archive/Resurrection/
                                        pdf/res24.pdf.

Wilkes:2000:SMD

[Wil00c] M. V. Wilkes. Slave memories and dynamic storage allocation.

In Hill et al. [HJS00], pages 371–

372. ISBN 1-55860-539-8. LCCN QA76.9.A73 H55 2000. URL

http://books.elsevier.com/
bookscat/links/details.asp?
isbn=1558605398;   http://
www.loc.gov/catdir/description/
els033/99044480.html;   http:
//www.loc.gov/catdir/toc/
els033/99044480.html;   https:
//archive.org/details/readingsincomput0000hill/
page/n9/mode/2up?q=Slave+
memories;   https://shop.
elsevier.com/books/readings-
in-computer-architecture/
hill/978-0-08-057364-9.

Wilkes:2000:CQD

[Wil00d] Maurice Wilkes. Comments, queries, and debates: Was Babbage caught in the act? IEEE Annals of the History of Computing, 22(4):69–70, October/December 2000. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2000/pdf/a4069.
pdf.

Wilkes:2000:IBA

[Wil00e]  Maurice V. Wilkes.  Intro-

duction to “Babbage’s Analytical Engine Plans 28 and 28a — The Programmer’s Interface” by Allan Bromley. IEEE Annals of the History of Computing, 22(4):4–5, October/December 2000. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://dlib.computer.
org/an/books/an2000/pdf/a4004.
pdf;   http://ieeexplore.

                                        ieee.org/iel5/85/19196/00887985.
                                        pdf;   http://www.computer.
                                        org/annals/an2000/a4004abs.
                                        htm.

Wilkes:2001:HPM

[Wil01a]  M. V. Wilkes.  High per-

formance memory systems. IEEE Transactions on Computers, 50(11):1105, November 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=966488.

Wilkes:2001:JPLa

[Wil01b] M. V. Wilkes. John Pinkerton and Lyons Electronic Office. Computing and Control Engineering Journal, 12(3):139–144, June 2001. CODEN CCEJEL. ISSN 0956-3385 (print), 1741-0460 (electronic). URL https:

                                        //digital-library.theiet.
                                        org/content/journals/10.1049/
                                        cce_20010308.

Wilkes:2001:JPLb

[Wil01c] M. V. Wilkes. John Pinkerton and Lyons Electronic Office. Engineering Science & Education Journal, 10(5):183–188, October 2001. CODEN ESEJEJ. ISSN 0963-7346 (print), 2051-2147 (electronic). URL https:

                                        //digital-library.theiet.
                                        org/content/journals/10.1049/
                                        esej_20010506.

Wilkes:2001:MGF

[Wil01d] Maurice V. Wilkes. The memory gap and the future of high performance memories. ACM SIGARCH Computer Architecture News, 29(1):2–7, March 2001. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Wilkes:2001:F

[Wil01e] Sir Maurice Wilkes FREng FRS.

Foreword. In Schonland: Scientist and Soldier [Aus01], page ?? ISBN 0-7503-0501-0 (hardcover), 1-4200-3357-3 (PDF ebook). LCCN QC16.S2646 A97 2001; Q143.S36 A87 2001. URL

http://www.crcnetbase.com/
isbn/9781420033571.

Wilkes:2002:PRN

[Wil02a] Maurice Wilkes. A personal revisitation of neural nets. Web document, November 1,

2002.  URL https://www.
cl.cam.ac.uk/archive/mvw1/
neural-nets-text-tex.pdf.

Wilkes:2002:CBH

[Wil02b] Maurice V. Wilkes. Charles Babbage and his world. Notes and Records of the Royal Society of London, 56(3):353–365, September 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

URL http://www.jstor.org/
stable/3557738.

Wilkes:2003:LUA

[Wil03]  Maurice V. Wilkes.  Let-

ters: The Use and Abuse of SPEC: An ISCA Panel, IEEE Micro, July-Aug. 2003, pp. 73–77. IEEE Micro, 23(6):5, November/December 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/mi/2003/06/m6005.pdf;
                                         http://csdl.computer.org/
                                        dl/mags/mi/2003/06/m6005.
                                        htm.

Wilkes:2004:CSA

[Wil04a] Maurice Wilkes. Computers for specialized application areas. In Herbert and Jones [HJ04], pages 317–318. ISBN 0-387-20170-X. LCCN QA76.9.S88 H45 2004.

Wilkes:2004:PCP

[Wil04b] Maurice Wilkes. Progress in computers: Prestige lecture delivered to IEE, Cambridge. Web document, February 5,

                                        2004.  URL https://www.
                                        cl.cam.ac.uk/archive/mvw1/
                                        Progress_in_Computers_IEE_
                                        Cambridge_2004_web.pdf.

Wilkes:2005:REV

[Wil05] Maurice Wilkes. Recollections of early vacuum tube circuits. In Copeland [Cop05], page ?? ISBN 0-19-856593-3 (hardcover). LCCN QA75 .A43 2005.

                                        URL http://ukcatalogue.
                                        oup.com/product/9780198565932.
                                        do;   http://www.oxfordscholarship.
                                        com/oso/public/content/maths/
                                        9780198565932/toc.html.

Wilkes:2006:WRE

[Wil06] Maurice Wilkes. What I remember of the ENIAC. IEEE Annals of the History of Computing, 28 (2):30–31, April/June 2006. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Wilkes:2012:REV

[Wil12] Maurice Wilkes. Recollections of early vacuum tube circuits. In Copeland [Cop12], chapter 18, pages 345–348. ISBN 0-19-856593-3 (hardcover), 0-19-960915-2 (paperback), 0-19-164434-X (e-book). LCCN QA75 .C638 2012. UK £14.99.

Wilby:2013:FGT

[Wil13]  Dave Wilby.  ‘Flash gor-

don’ tech: How Sir Maurice Wilkes made practical computers possible: ENIAC — a time before integrated circuits. The Register, ??(??):??, June 26, 2013. URL https://

www.theregister.com/Print/
2013/06/26/maurice_wilkes_
centenery/.

Wilkes:2021:BWD

[Wil21] Maurice Wilkes. The best way to design an automatic calculating machine (1951). In Lewis [Lew21], pages 165–168. ISBN 0-262-04530-3. LCCN QA76 .I34 2021. Reprinted from [Wil81a].

Williams:1951:MUC

[WK51] F. C. Williams and T. Kilburn, editors. Manchester University computer: inaugural conference [held at the University on the 9th, 10th, 11th and 12th July, 1951]. Tillotsons, Bolton, England, July 1951.

LCCN ???? URL https:

//ieeemilestones.ethw.org/
w/images/5/5d/Inaug_Conf_
Ref_13.pdf.

Wilkes:2003:TKC

[WK03] Maurice Wilkes FRENG FRS and Hilary J. Kahn. Tom Kilburn CBE FREng. 11 August 1921–17 January 2001. Biographical Memoirs of Fellows of the Royal Society, 49:283–297, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:

//royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.
2003.0016.

Wilkes:1986:RMC

[WLC86] Maurice V. Wilkes, J. A. N. Lee, and Paul E. Ceruzzi. Reviews: M. Campbell-Kelly and M. R. Williams (eds.), The Moore School Lectures; Computer Museum Slide Set; K. Zuse. Der Computer, Mein Lebenswerk; capsule reviews. Annals of the History of Computing, 8(3):315–319, July/ September 1986. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/
an/books/an1986/pdf/a3315.
pdf;   http://www.computer.
org/annals/an1986/a3315abs.
htm.

Wilkes:1972:ALA

[WM72] M. V. Wilkes and W. D. Manville. The associative language AL-1. International Journal of Computer and Information Sciences, 1(4):315–327, December 1972. CODEN IJCIAH. ISSN 0091-7036.

Wilkes:1968:DMA

[WN68] M. V. Wilkes and R. M. Needham. The design of multipleaccess computer systems: part

2. The Computer Journal, 10(4):315–320, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/10/4/315.full.pdf+
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_04/100315.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
10/Issue_04/tiff/315.tif;
 http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/316.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/317.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/318.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/319.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_
04/tiff/320.tif.

Wilkes:1979:CCC

[WN79] M. V. Wilkes and R. M. Needham. The Cambridge CAP Computer and Its Operating System, volume 6 of The computer science library; operating and programming systems series. North-Holland, Amsterdam, The Netherlands, 1979. ISBN 0-444-00357-6, 0-444-00358-4. x + 165 pp. LCCN QA76.8.C16 W54.

Wilkes:1980:CMD

[WN80] Maurice V. Wilkes and Roger M. Needham. The Cambridge Model Distributed System. Operating Systems Review, 14(1): 21–29, January 1980. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Worsley:1950:ED

[Wor50] B. H. Worsley. The EDSAC demonstration. In Report of a conference on High Speed Automatic Calculating Machines. University Mathematical Laboratory, Cambridge University, Cambridge, UK, January 1950. Reprinted in [Ran82, pages 423– 429].

Worsley:1953:SPR

[Wor53] Beatrice Helen Worsley. Serial Programming for Real and Idealised Digital Calculating Machines. Ph.D. thesis, University of Cambridge, Cambridge, UK, 1953. URL https:/

                                        /www.mathgenealogy.org/id.
                                        php?id=42178.

Wilkes:1954:EMC

[WPB54] M. V. Wilkes, M. Phister, and S. A. Barton. Experience with marginal checking and automatic routining of the EDSAC. In ????, editor, Automatic digital computation. [Proceedings of a symposium held at the National Physical Laboratory on March 25, 26, 27 & 28, 1953], pages 239–?? HMSO, London, UK, 1954. LCCN QA76 .T4 1953.

Wilkes:1989:EMC

[WPB89] M. V. Wilkes, M. Phister, and S. A. Barton. Experience with marginal checking and automatic routing of the EDSAC. In Williams and Campbell-Kelly [WCK89], pages 446–453. ISBN 0-262-23136-0. LCCN QA76.17 .E171 1989.

Wilkes:1947:TPR

[WR47] M. V. Wilkes and J. A. Ramsay. A theory of the performance of radar on ship targets. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 43(??):220–231, April 1947. CODEN PCPSA4. ISSN 0008-1981. URL https:

//www.cambridge.org/core/
journals/mathematical-proceedings-
of-the-cambridge-philosophical-
society/article/abs/theory-
of-the-performance-of-radar-
on-ship-targets/F9063FA219E8A34B073F74DB08C41A3A.

Wilkes:1948:UMU

[WR48] Maurice V. Wilkes and William Renwick. An ultrasonic memory unit for the EDSAC. Electronic Engineering, 20(245): 208–213, July 1948. CODEN ELEGAP. ISSN 0013-4902.

URL http://www.cs.man.ac.
uk/CCS/Archive/misc/EDSAC/
EDSAC%20Memory%20Text.pdf.

Wilkes:1949:EEC

[WR49] Maurice V. Wilkes and William Renwick. The EDSAC — an electronic calculating machine. Journal of Scientific Instruments, 26(12):385–391, December 1949. CODEN JSINAY. ISSN 0368-4253, 0950-7671.

                                        URL https://iopscience.
                                        iop.org/article/10.1088/0950-
                                        7671/26/12/301. Reprinted in

[Ran82, pp. 417–422].

Wilkes:1950:E

[WR50a] M. V. Wilkes and W. Renwick. The EDSAC. In Report of a Conference on High Speed Automatic Calculating Machines, 22–25 June 1949, pages 9–11. University Mathematical Laboratory, Cambridge, UK, January 1950. Reprinted in [Ran82, §8.5].

Wilkes:1950:EED

[WR50b] M. V. Wilkes and W. Renwick. The EDSAC (Electronic Delay Storage Automatic Calculator). Mathematical Tables and Other Aids to Computation, 4(30):61– 65 (1 plate), April 1950. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Wilkes:1973:E

[WR73] M. V. Wilkes and W. Renwick. The EDSAC. In Randell [Ran73], pages 417–421. ISBN 0-387-06169-X (New York), 3-540-06169-X. LCCN TK7888.3 .R36.

Wilkes:1975:SPE

[WR75] M. V. Wilkes and W. Renwick. Stored program electronic computers. In Randell [Ran75], pages 389–393. ISBN 0-387-07114-8, 3-540-07114-8, 3-642-96244-0, 3-642-96242-4 (ebook). LCCN ?TK7888.3 .R36 1975.

Wilkes:1982:E

[WR82] M. V. Wilkes and W. Renwick. The EDSAC (1949). In Randell [Ran82], pages 417– 421. ISBN 0-387-11319-3, 3-540-11319-3. LCCN TK7885.A5 O741 1982.

Wilkes:1958:CCU

[WRW58a] M. V. Wilkes, W. Renwick, and D. J. Wheeler. Computer control unit. Journal of the Institution of Electrical Engineers, 4(37):24–25, January 1958. CODEN JISEAL. ISSN 0099-2887 (print), 2054-0612 (electronic). URL https:

//digital-library.theiet.
org/content/journals/10.1049/
jiee-3.1958.0009.

Wilkes:1958:DCU

[WRW58b] M. V. Wilkes, W. Renwick, and D. J. Wheeler. The design of the control unit of an electronic digital computer. Proceedings of the IEE — Part B: Radio and Electronic Engineering, 105(20):121–128, March 1958. CODEN ???? ISSN 2050-2656 (print), 2054-0434 (electronic). URL https:

//digital-library.theiet.

                                        org/content/journals/10.1049/
                                        pi-b-1.1958.0267.

Wilkes:1953:MPD

[WS53] M. V. Wilkes and J. B. Stringer. Micro-programming and the design of the control circuits in an electronic digital computer. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 49(2):230–238, April 1953. CODEN PCPSA4. ISSN 0008-1981. URL https:

                                        //www.cambridge.org/core/
                                        journals/mathematical-proceedings-
                                        of-the-cambridge-philosophical-
                                        society/article/abs/microprogramming-
                                        and-the-design-of-the-control-
                                        circuits-in-an-electronic-
                                        digital-computer/4F7EDCBEB735637790E7170503908323.

Wilkes:1971:MDC

[WS71] M. V. Wilkes and J. B. Stringer. Microprogramming and the design of the control circuits in an electronic computer. In Bell and Newell [BN71], chapter 28, pages 335–340. ISBN 0-07-004357-4. LCCN TK7888.3 .B4.

Wilkes:1982:MDC

[WS82] M. V. Wilkes and J. B. Stringer. Microprogramming and the design of the control circuits in an electronic digital computer. In Computer Structures: Principles and Examples [BSN82], pages 158–163. ISBN 0-07-057302-6. LCCN QA76.9.A73 C65; TK7888.3 .B37 1982.

Wilkes:1986:RKW

[WS86] M. V. Wilkes and K. W. Smillie. Reviews: K. L. Wildes and N. Lindgren, A Century of Electrical Engineering and Computer Science at MIT; J. G. Kemeny and T. E. Kurtz, Back to BASIC; capsule reviews. Annals of the History of Computing, 8(1):94, 96–98, January/ March 1986. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/
an/books/an1986/pdf/a1094.
pdf;   http://www.computer.
org/annals/an1986/a1094abs.
htm.

Wilkes:1988:WFS

[WU88] M. V. Wilkes and M. Underwood. The Winthrop family and St John’s College, Cambridge. The Eagle, ??(??):xxx–???, ???? 1988.

Weekes:1947:AOR

[WW47] K. Weekes and Maurice Vincent Wilkes. Atmospheric oscillations and the resonance theory. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 192(1028):80–99, December 23, 1947. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https:

//royalsocietypublishing.
org/doi/10.1098/rspa.1947.
0139.

Wilkes:1956:MTA

[WW56a] M. V. Wilkes and D. W.

Willis. A magnetic-tape auxiliary storage system for the ED-SAC. Proceedings of the IEE — Part B: Radio and Electronic Engineering, 103(2S):337–345, April 1956. CODEN PEEBAE. ISSN 0369-8890 (print), 2054-0418 (electronic). URL https:

//digital-library.theiet.
org/content/journals/10.1049/
pi-b-1.1956.0070.

Wilkes:1956:MTS

[WW56b] M. V. Wilkes and D. W. Willis. A magnetic-tape auxiliary storage system for the EDSAC. Proceedings of the IEE — Part B: Radio and Electronic Engineering, 103(2S): 337–345, April 1956. ISSN 2050-2656 (print), 2054-0434 (electronic). URL https:

//digital-library.theiet.
org/content/journals/10.1049/
pi-b-1.1956.0070.

Wilkes:1957:NMC

[WW57] M. V. Wilkes and D. J. Wheeler. Note on “A method for computing certain inverse functions” (in Technical Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 11(59):204, July 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL https://www.
jstor.org/stable/2002084.

Wilkes:1958:ASM

[WW58] M. V. Wilkes and D. J. Wheeler. Auxiliary storage on magnetic tape in EDSAC 2. In ????, editor, Congresso Internacional de Automática Madrid [13– 18 October 1958], pages 185– ?? ????, ????, 1958. LCCN

???? URL http://www.fdi.
ucm.es/migs/multimedia/01-
CongresoAutomatica/#/3 [photo
gallery].

Wilkes:1959:DHS

[WW59] M. V. Wilkes and D. J. Wheeler.

The design of high-speed photoelectric tape readers. Proceedings of the IEE — Part B: Radio and Electronic Engineering, 106(29):450, September 1959. CODEN PEEBAE. ISSN 0369-8890 (print), 2054-0418 (electronic). URL https:

//digital-library.theiet.
org/content/journals/10.1049/
pi-b-2.1959.0293.

Wilkes:1979:CDC

[WW79] M. V. Wilkes and D. J. Wheeler.

The Cambridge digital communications ring. In R. Rosenthal and N. B. Meisner, editors, Proceedings of the Local Area Communications Network Symposium: Boston, May 7–9, 1979, pages 47–62. National Technical Information Service, Springfield, VA, USA, 1979. LCCN ????

Wilkes:1951:PPE

[WWG51] Maurice V. Wilkes, F.R.S., David J. Wheeler, and Stanley Gill. The Preparation of Programs for an Electronic Digital Computer. Addison-Wesley, Reading, MA, USA, 1951. x + 170 pp. LCCN QA76.5 .W55 1951. With Special Reference to the EDSAC and the Use of a Library of Subroutines. With a foreword by D. R. Hartree. See also second edition [WWG57], and reprint [WWG82].

Wilkes:1954:IPA

[WWG54] Maurice V. Wilkes, D. Wheeler, and Stanley Gill. Introduction to Programming for an Automatic Digital Calculating Machine and Users’ Guide to the EDSAC: a Supplement to The Preparation of Programs for an Electronic Digital Computer. Cambridge University Press, Cambridge, UK, 1954. Part I by D. R. Hartree. Part II edited by E. N. Mutch.

Wilkes:1957:PPE

[WWG57] Maurice V. Wilkes, David J. Wheeler, and Stanley Gill. The Preparation of Programs for an Electronic Digital Computer. Addison-Wesley Mathematics Series. Addison-Wes-ley, Reading, MA, USA, second edition, 1957. xiv + 238 pp. LCCN QA76.5 .W52 1957.

URL https://archive.org/
details/programsforelect00wilk/
;   https://b-ok.org/book/
3668116/b363ff;.  Preface

dated August 1957. See also first edition [WWG51] and review [Cor58].

Wilkes:1982:PPE

[WWG82] M. V. (Maurice Vincent) Wilkes, David J. Wheeler, and Stanley Gill. The Preparation of Programs for an Electronic Digital Computer: with Special Reference to the EDSAC and the Use of a Library of Subroutines, volume 1 of Charles Babbage Institute reprint series for the history of computing. Tomash Publishers, Los Angeles, CA, USA, 1982. ISBN 0-262-23118-2 (MIT Press 1984), 0-938228-03-X. xxxi + 167 pp. LCCN QA76.6 .W545 1982. With a new introduction by Martin Campbell-Kelly.